### 20 Mar 2026 v2

# Congruences between Klingen-Eisenstein series and cusp forms on $\mathrm{U}_{n,n}$

Abstract.
In this paper, we study congruences of Hecke eigenvalues between Hermitian Klingen–Eisenstein series and cusp forms on the unitary group $\mathrm{U}_{n,n}$ defined over the rational number field $\mathbb{Q}$. We also prove the rationality of the space of Hermitian automorphic forms and the integrality of their Hecke eigenvalues.
Key words and phrases:
Hermitian modular form, Unitary groups, Congruences
2020 Mathematics Subject Classification:
11F33, 11F55, 11F30, 11F67
This work was supported by JST SPRING, Grant Number JPMJSP2110.
1. Introduction
Studying congruences between Eisenstein series and cusp forms is fundamental to the arithmetic theory of automorphic forms. These congruences are not merely arithmetic curiosities but are deeply intertwined with the special values of $L$-functions and have far-reaching applications in Iwasawa theory. The existence of a congruence often implies that the $p$-adic properties of an $L$-value control the arithmetic of the associated Galois representations, a philosophy central to the proof of the Iwasawa Main Conjecture.
A foundational instance of this principle is the work of Skinner and Urban [32], who established the Main Conjecture for $\mathrm{GL}_{2}$. Their strategy involved constructing congruences between Eisenstein series and cusp forms on the unitary similitude group $\mathrm{GU}(2,2)$.
Analytic methods for constructing congruences using the pullback formula have been extensively developed, primarily in the context of symplectic groups $\mathrm{Sp}(n)$. Recently, Katsurada [20] and Katsurada-Mizumoto [19] established a refined criterion for the existence of congruences between Klingen-Eisenstein series and cusp forms on $\mathrm{Sp}(n)$. Their method utilizes the pullback formula to explicitly relate the Petersson inner product of a cusp form and a restricted Eisenstein series to special values of standard $L$-functions.
In this paper, we extend these powerful techniques to the case of (vector-valued) Hermitian automorphic forms on the unitary group $\mathrm{U}_{n,n}$ defined over the rational number field $\mathbb{Q}$. The primary objective of this work is to establish a precise criterion for the existence of congruences between Hermitian Klingen-Eisenstein series and Hecke cusp forms on $\mathrm{U}_{n,n}$.
Specifically, let $r<n$ be a positive integer. We consider a Hecke cusp form $f$ on the smaller unitary group $\mathrm{U}_{r,r}$ and its associated Klingen-Eisenstein series $f]_{r}^{n}$ (defined in Section 4.2) on $\mathrm{U}_{n,n}$. The main result of this paper (Theorem 6.2) asserts that if the algebraic part of the special value of the standard $L$-function associated with $f$, denoted by $L(s,f,\mathrm{St})$, is divisible by a prime ideal $\mathfrak{p}$ (under appropriate normalization and conditions on the prime $p$), then there exists a Hecke cusp form $F$ on $\mathrm{U}_{n,n}$ such that
$F\equiv_{ev}[f]_{\mu}^{n}\pmod{\mathfrak{P}}$,
where $\mathfrak{P}$ is a prime ideal of the Hecke field lying above $\mathfrak{p}$, and $\equiv_{ev}$ denotes the congruence of eigenvalues for all Hecke operators in the spherical Hecke algebra discussed in Section 3.
To prove this theorem, we first utilize the vector-valued differential operators $\mathbb{D}_{\bm{k},\bm{l}}$ constructed in our previous work [34], which are designed to ensure that the pullback of an automorphic form remains automorphic on the subgroup. By applying these operators to the Eisenstein series on $\mathrm{U}_{n,n}$ and restricting it to $\mathrm{U}_{n_{1},n_{1}}\times\mathrm{U}_{n_{2},n_{2}}$, we derive a pullback formula (Theorem 4.8). At the same time, we establish the necessary arithmetic framework by defining an integral structure on the space of automorphic forms and proving the rationality of the space of Hermitian automorphic forms (Theorem 5.12).
Finally, we apply the strategy of Katsurada-Mizumoto [19] to the Hermitian setting by analyzing the denominators of the Fourier coefficients of the Klingen-Eisenstein series $f]_{r}^{n}$.
We now describe our main result (Theorem 6.2) in more detail.
Let $n$ be a positive integer and let $\mu$ be a positive even integer with $\mu\geq 2n$. Let $(\bm{k},\bm{l})$ be a pair of dominant integral weights whose lengths satisfy $\ell(\bm{k}),\ell(\bm{l})\leq n$. Let $f$ be a Hecke eigen cusp form on the unitary group $\mathrm{U}_{r,r}$ of signature $(r,r)$, and denote by $f]_{\mu}^{n}$ the Klingen-Eisenstein series on $\mathrm{U}_{n,n}$. We write $\mathbb{Q}(f)$ for the Hecke field generated by the eigenvalues of $f$.
To formulate the congruence criterion, we introduce a constant
$\mathcal{C}_{n,\mu}(f)=\frac{\mathbb{L}_{F}(n,\mu)}{\mathbb{L}_{F}(2r,\mu)}\,\mathbb{L}\bigl((\mu+1)/2-r,f\bigr).$
(See Section 5 for the precise definitions and normalization.)
Let $\mathfrak{p}$ be a prime ideal of $\mathbb{Q}(f)$. Assume that $\mathfrak{p}$ satisfies
$v_{\mathfrak{p}}\!\left(\overline{\mathcal{C}_{2n,\mu}(f)}\,A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})\,\overline{A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})}\right)<0,$
where $v_{\mathfrak{p}}$ denotes the $\mathfrak{p}$-adic valuation, and $A_{[f]_{\mu}^{n}}(\ast,S_{0})$ is an $S_{0}$-th Fourier coefficient of $f]_{\mu}^{n}$ defined as (2.6). We further impose several natural technical conditions on $\mathfrak{p}$ ensuring integrality.
Then there exists a Hecke eigen cusp form $F$ on $\mathrm{U}_{n,n}$ such that
$F\equiv_{ev}[f]_{\mu}^{n}\pmod{\mathfrak{P}}$
for some prime ideal $\mathfrak{P}$ of the Hecke field of $F$ lying above $\mathfrak{p}$. Here the notation $F\equiv_{ev}G\pmod{\mathfrak{P}}$ means that the Hecke eigenvalues of $F$ and $G$ are congruent modulo $\mathfrak{P}$ for all Hecke operators in the spherical Hecke algebra.
Moreover, under a stronger condition on the $\mathfrak{p}$-adic valuation of the above algebraic quantity, we obtain a refined congruence modulo a higher power of $\mathfrak{P}$.
This paper is organized as follows: Section 2 provides the preliminary definitions of Hermitian automorphic forms on $\mathrm{U}_{n,n}$ and their Fourier expansions. Section 3 is devoted to the local theory of Hecke algebras, describing the structure at inert, ramified, and split places. Section 4 introduces the differential operators and states the pullback formula (Theorem 4.8). Section 5 discusses the arithmetic properties and rationality of the space of automorphic forms (Theorem 5.12). Section 6 contains the proof of the main result. Section 7 provides explicit numerical examples for $n=2$ and $n=3$.
Acknowledgements
The author is grateful to T. Ikeda for his guidance and support as the author’s doctoral advisor, and to H. Katsurada for valuable comments. This work was supported by the Japan Science and Technology Agency (JST) SPRING Program, Grant Number JPMJSP2110.
Notation
We denote by $\mathrm{M}_{m,n}(R)$ the set of $m\times n$ matrices with entries in $R$. In particular, we put $\mathrm{M}_{n}(R):=\mathrm{M}_{n,n}(R)$. Let $I_{n}$ be the identity element of $\mathrm{M}_{n}(R)$ and $e_{ij}$ the matrix with $1$ at the $(i,j)$-th entry and 0 elsewhere. Let $\det(X)$ be the determinant of $X$ and $\operatorname{Tr}(X)$ the trace of $X$, $^{t}X$ the transpose of $X$ for a square matrix $x\in\mathrm{M}_{n}(R)$. Let $\mathrm{GL}_{n}(R)\subset\mathrm{M}_{n}(R)$ be a general linear group of degree $n$.
Let $K$ be a quadratic extension field of $K_{0}$ with the non-trivial automorphism $\rho$ of $K$ over $K_{0}$, and we often put $\overline{x}=\rho(x)$ for $k\in K$. We put $\overline{X}=(\overline{x_{ij}})$ and ${X^{*}}={}^{t}\overline{X}$ for $X=(x_{ij})\in\mathrm{M}_{m,n}(K)$. Let $B^{*}$} be the transpose of $\overline{B}$ and $\overline{B}$ the complex conjugate of $B$.
Let $K$ be an algebraic number field, and let $\mathfrak{p}$ be a prime ideal of $K$. We denote by $K_{\mathfrak{p}}$ a $\mathfrak{p}$-adic completion of $K$
and by $\mathcal{O}_{K}$ (resp. $\mathcal{O}_{\mathfrak{p}}$) the integer ring of $K$ (resp. $K_{\mathfrak{p}}$). Let $v_{\mathfrak{p}}$ be the additive valuation of $K_{\mathfrak{p}}$ normalized by $v_{\mathfrak{p}}(\varpi_{\mathfrak{p}})=1$, where $\varpi_{\mathfrak{p}}$ is a uniformizer of $K_{\mathfrak{p}}$. We put $\mathcal{O}_{(\mathfrak{p})}=K\cap\mathcal{O}_{\mathfrak{p}}$. Let $\mathrm{Her}_{n}(\mathbb{C})\subset\mathrm{M}_{n}(\mathbb{C})$ be the set of Hermitian matrices. For an element $X\in\mathrm{Her}_{n}(\mathbb{C})$, we denote by $X>0$ (resp. $X\geq 0$) that $X$ is a positive definite matrix (resp. a non-negative definite matrix). For a subset $S\subset\mathrm{Her}_{n}(\mathbb{C})$, we denote by $S_{>0}$ (resp. $S_{\geq 0}$) the subset of positive definite (resp. non-negative definite) matrices in $S$. If a group $G$ acts on a set $V$, then we denote by $V^{G}$ the $G$-invariant subspace of $V$.
Let ${\det}^{k}$ be the one-dimensional representation of $\mathrm{GL}_{n}(\mathbb{C})$ given by the $k$-th power of the determinant, and let $\mathrm{Sym}^{l}$ be the $l$-th symmetric power representation of $\mathrm{GL}_{n}(\mathbb{C})$. For a representation $(\rho,V)$, we denote by $(\rho^{*},V^{*})$ the contragredient representation of $(\rho,V)$.
We denote by $(x)^{(r)}=x(x+1)\cdots(x+r-1)$ (resp. $(x)_{(r)}=x(x-1)\cdots(x-r+1)$) the ascending (resp. descending) Pochhammer symbol.
2. Hermitian automorphic forms
Let $K$ be a quadratic imaginary extension of a totally real field $F$, and let $\mathbf{a}$ (resp. $\mathbf{h}$) denote the set of infinite (resp. finite) places of $F$. We fix a CM type $\Sigma_{K}$ for $K$ (i.e. $\Sigma_{K}$ contains a choice of exactly one representative from each pair of complex conjugate embeddings of $K$). We often canonically identify CM type $\Sigma_{K}$ with $\mathbf{a}$, and we denote by $\sigma_{v}$ the embedding $K\hookrightarrow\mathbb{C}$ corresponding to $v\in\mathbf{a}$.
We put $m=m_{F}:=\#\mathbf{a}=[F:\mathbb{Q}]$. We put $K_{v}=\prod_{w|v}K_{w}$ and $\mathcal{O}_{K_{v}}=\prod_{w|v}\mathcal{O}_{K_{w}}$ for a place $v$ of $F$. Let $\mathbb{A}=\mathbb{A}{F}$ be the adele ring of $F$, and $\mathbb{A}_{0}$, $\mathbb{A}_{\infty}$ the finite and infinite parts of $\mathbb{A}$, respectively.
We put $J_{n}=\begin{pmatrix}O_{n}&I_{n}\\ -I_{n}&O_{n}\\ \end{pmatrix}$. The unitary group $\mathrm{U}_{n,n}$ and the unitary similitude group $\mathrm{GU}_{n,n}$ are algebraic groups defined over the totally real field $F$. For any $F$-algebra $R$, their $R$-points are given by
$\mathrm{U}_{n,n}(R)$ $=\{g\in\mathrm{GL}_{2n}(K\otimes_{F}R)\mid{g^{*}}J_{n}g=J_{n}\},$
$\mathrm{GU}_{n,n}(R)$ $=\{g\in\mathrm{GL}_{2n}(K\otimes_{F}R)\mid{g^{*}}J_{n}g=\nu(g)J_{n},\ \nu(g)\in R^{\times}\}.$
We call $\nu(g)$ the similitude factor of $g$. If we put
$\pi(g)=\begin{pmatrix}\nu(g)I_{n}&0_{n}\\ 0_{n}&I_{n}\end{pmatrix}^{\!-1}\hskip-5.0ptg\in\mathrm{U}_{n,n}$
for $g\in\mathrm{GU}_{n,n}$, we have the isomorphism
$\mathrm{GU}_{n,n}$ $\cong$ $\mathrm{U}_{n,n}\rtimes\mathbb{G}_{m},$
$g$ $\mapsto$ $\left(\pi(g),\ \nu(g)\right),.$ (2.1)
We also define other unitary groups $\mathrm{U}(n,n)$, $\mathrm{GU}(n,n)$, $\mathrm{GU}^{+}(n,n)$, and $\mathrm{U}(n)$ by
$\mathrm{U}(n)$ $=\{g\in\mathrm{GL}_{n}(\mathbb{C})\mid{g^{*}}g=I_{n}\},$
$\mathrm{U}(n,n)$ $=\{g\in\mathrm{GL}_{2n}(\mathbb{C})\mid{g^{*}}J_{n}g=J_{n}\},$
$\mathrm{GU}(n,n)$ $=\{g\in\mathrm{GL}_{2n}(\mathbb{C})\mid{g^{*}}J_{n}g=\nu(g)J_{n},\ \nu(g)\in\mathbb{R}^{\times}\},$
$\mathrm{GU}^{+}(n,n)$ $=\{g\in\mathrm{GL}_{2n}(\mathbb{C})\mid{g^{*}}J_{n}g=\nu(g)J_{n},\ \nu(g)\in\mathbb{R}^{\times}_{>0}\}.$
We put $\widetilde{G}_{n}=\mathrm{GU}_{n,n}(F)$ and
$\widetilde{G}^{+}_{n}=\{g\in\mathrm{GU}_{n,n}(F)\mid\nu(g)\in F_{>0}\},$
where
$F_{>0}=\{x\in F\mid\sigma(x)>0\text{ for all embeddings }\sigma:F\hookrightarrow\mathbb{R}\}.$
Let $\widetilde{G}_{n,v}=\mathrm{GU}_{n,n}(F_{v})$ for a place $v$ of $F$,
$\widetilde{G}_{n}(\mathbb{A})=\mathrm{GU}_{n,n}(\mathbb{A}),\quad\widetilde{G}_{n,0}=\mathrm{GU}_{n,n}(\mathbb{A}_{0}),$
$\widetilde{G}_{n,\infty}=\mathrm{GU}_{n,n}(F\otimes_{\mathbb{Q}}\mathbb{R})=\prod_{v\in\mathbf{a}}\widetilde{G}_{n,v}=\prod_{v\in\mathbf{a}}\mathrm{GU}(n,n).$
We also set
$\widetilde{G}^{+}_{n,\infty}=\prod_{v\in\mathbf{a}}\mathrm{GU}^{+}(n,n),\qquad\widetilde{G}^{+}_{n}(\mathbb{A})=\prod_{v\in\mathbf{h}}\mathrm{GU}_{n,n}(F_{v})\times\widetilde{G}^{+}_{n,\infty}$.
Similarly, for the unitary group $\mathrm{U}_{n,n}$, we define
$G_{n}=\mathrm{U}_{n,n}(F),\quad G_{n,v}=\mathrm{U}_{n,n}(F_{v}),\quad\text{etc.}$
We define $\mathfrak{K}_{n,v}$ and $\widetilde{\mathfrak{K}}_{n,v}$ by
$\mathfrak{K}_{n,v}=\left\{\begin{array}[]{ll}\mathrm{U}_{n,n}(\mathcal{O}_{F_{v}})&(v\in\mathbf{h}),\\ \mathrm{U}(n)\times\mathrm{U}(n)&(v\in\mathbf{a}),\end{array}\right.\quad\text{and}\quad\widetilde{\mathfrak{K}}_{n,v}=\left\{\begin{array}[]{ll}\mathrm{GU}_{n,n}(\mathcal{O}_{F_{v}})&(v\in\mathbf{h}),\\ \mathrm{U}(n)\times\mathrm{U}(n)&(v\in\mathbf{a}).\end{array}\right$.
Then $\mathfrak{K}_{n,v}$ (resp. $\widetilde{\mathfrak{K}}_{n,v}$) is isomorphic to a maximal compact subgroup of $G_{n,v}$ (resp. $\widetilde{G}^{+}_{n,v}$) for each place $v\in\mathbf{a}$. We fix a maximal compact subgroup of $G_{n,v}$ (resp. $\widetilde{G}_{n,v}$), which is also denoted by $\mathfrak{K}_{n,v}$ (resp. $\widetilde{\mathfrak{K}}_{n,v}$) by abuse of notation. We put $\mathfrak{K}_{n,0}=\prod_{v\in\mathbf{h}}\mathfrak{K}_{n,v}$, $\mathfrak{K}_{n,\infty}=\prod_{v\in\mathbf{a}}\mathfrak{K}_{n,v}$, $\widetilde{\mathfrak{K}}_{n,0}=\prod_{v\in\mathbf{h}}\widetilde{\mathfrak{K}}_{n,v}$ and $\widetilde{\mathfrak{K}}_{n,\infty}=\prod_{v\in\mathbf{a}}\widetilde{\mathfrak{K}}_{n,v}$.
2.1. As analytic functions on Hermitian symmetric spaces
We have the identification
$\mathrm{M}_{n}(\mathbb{C})$ $\cong\mathrm{Her}_{n}(\mathbb{C})\otimes_{\mathbb{R}}\mathbb{C},$
$Z$ $\mapsto\mathrm{Re}(Z)+\sqrt{-1}\,\mathrm{Im}(Z),$
where the Hermitian real and imaginary parts are given by
$\mathrm{Re}(Z)$ $=\tfrac{1}{2}(Z+{Z^{*}}),$
$\mathrm{Im}(Z)$ $=\tfrac{1}{2\sqrt{-1}}(Z-{Z^{*}}).$
Let $\mathfrak{H}_{n}$ denote the Hermitian upper half space of degree $n$:
$\mathfrak{H}_{n}=\{Z\in\mathrm{M}_{n}(\mathbb{C})\mid\mathrm{Im}(Z)>0\}.$
The group $\widetilde{G}^{+}_{n,\infty}=\prod_{v\in\mathbf{a}}\mathrm{GU}^{+}(n,n)$ acts on $\mathfrak{H}_{n}^{\mathbf{a}}$ by
$gZ=\bigl((A_{v}Z_{v}+B_{v})(C_{v}Z_{v}+D_{v})^{-1}\bigr)_{v\in\mathbf{a}},$
for $g=\begin{pmatrix}A_{v}&B_{v}\\ C_{v}&D_{v}\end{pmatrix}_{v\in\mathbf{a}}\in\widetilde{G}^{+}_{n,\infty}$ and $Z=(Z_{v})_{v\in\mathbf{a}}\in\mathfrak{H}_{n}^{\mathbf{a}}$. Set $\bm{i}_{n}:=(\sqrt{-1}I_{n})_{v\in\mathbf{a}}\in\mathfrak{H}_{n}^{\mathbf{a}}$.
For $g=(g_{v})_{v\in\mathbf{a}}=\begin{pmatrix}A_{v}&B_{v}\\ C_{v}&D_{v}\end{pmatrix}_{v\in\mathbf{a}}\in\widetilde{G}^{+}_{n,\infty}$ and $Z=(Z_{v})_{v\in\mathbf{a}}\in\mathfrak{H}_{n}^{\mathbf{a}}$, define
$\lambda(g,Z)=\bigl(\nu(g_{v})^{-1/2}(C_{v}Z_{v}+D_{v})\bigr)_{v\in\mathbf{a}},\quad\mu(g,Z)=\bigl(\nu(g_{v})^{-1/2}(\overline{C_{v}}{}^{t}Z_{v}+\overline{D_{v}})\bigr)_{v\in\mathbf{a}},$
and
$M(g,Z)=(\lambda(g,Z),\,\mu(g,Z)).$
For brevity, we write
$\lambda(g)=\lambda(g,\bm{i}_{n}),\quad\mu(g)=\mu(g,\bm{i}_{n}),\quad M(g)=M(g,\bm{i}_{n}).$
Let $(\rho,V)$ be an algebraic representation of $\widetilde{\mathfrak{K}}_{n,\infty}^{\mathbb{C}}:=\prod_{v\in\mathbf{a}}(\mathrm{GL}_{n}(\mathbb{C})\times\mathrm{GL}_{n}(\mathbb{C}))$ on a finite dimensional complex vector space $V$, and take a Hermitian inner product on $V$ such that
$\left<\rho(g)v,w\right>=\left<v,\rho({g^{*}})w\right$>
for any $g\in\widetilde{\mathfrak{K}}_{n,\infty}^{\mathbb{C}}$. Let $\omega:(\mathbb{R}_{>0})^{\mathbf{a}}\rightarrow\mathbb{C}^{\times}$ be a continuous character. For a $V$-valued function $F$ on $\mathfrak{H}_{n}^{\mathbf{a}}$, we put
$F|_{(\rho,\omega)}g(Z)=\omega(\nu(g))^{-1}\rho(M(g,Z))^{-1}F(g\langle Z\rangle)\quad(g\in\widetilde{G}^{+}_{n,\infty},\ Z\in\mathfrak{H}_{n}^{\mathbf{a}}).$
We set
$\widetilde{\Gamma}_{n}=\widetilde{G}^{+}_{n,\infty}\cap\mathrm{GL}_{2n}(\mathcal{O}_{K}).$
Definition 2.1.
We say that $F$ is a (holomorphic, level 1) Hermitian modular form of weight $(\rho,V)$ and character $\omega$ if $F$ is a holomorphic $V$-valued function on $\mathfrak{H}_{n}^{\mathbf{a}}$ and $F|_{(\rho,\omega)}g=F$ for all $g\in\widetilde{\Gamma}_{n}$. . (If $n=1$ and $F=\mathbb{Q}$, another holomorphy condition at the cusps is also needed.)
We denote by $M_{\rho}(\widetilde{\Gamma}_{n},\omega)$ a complex vector space of all Hermitian modular forms of weight $(\rho,V)$, and character $\omega$.
We set
$\Lambda_{n}:=\mathrm{Her}_{n}(K)\cap M_{n}(\mathfrak{d}_{K}^{-1}),$
where $\mathfrak{d}_{K}$ denotes the different ideal of $K$. We further denote by ${\Lambda_{n}}_{\geq 0}$ the subset of $\Lambda_{n}$ consisting of matrices that are non-negative definite for every embedding $\sigma\in\Sigma_{K}$.
Then, a modular form $F\in M_{\rho}(\widetilde{\Gamma}_{n},\omega)$ has the Fourier expansion
$F(Z)=\sum_{T\in{\Lambda_{n}}_{\geq 0}}a(F,T)\mathbf{e}\left(\sum_{v\in\mathbf{a}}(\operatorname{Tr}(T_{\sigma_{v}}Z_{v}))\right),$
where $a(F,T)\in V$, $\mathbf{e}(z)=\exp(2\pi\sqrt{-1}z)$. Here, $T_{\sigma_{v}}$ is the image of $T\in\mathrm{Her}_{n}(K)$ under the embedding corresponding to $\sigma_{v}\in\Sigma_{K}$. If $a(F,T)=0$ unless $T$ is positive definite,
we say that $F$ is a (holomorphic) Hermitian cusp form of weight $(\rho,V)$, and character $\omega$. We also denote by $S_{\rho}(\widetilde{\Gamma}_{n},\omega)$ a complex vector space of all cusp forms of weight $(\rho,V)$, and character $\omega$.
Write the variable $Z=(X_{v}+\sqrt{-1}Y_{v})_{v\in\mathbf{a}}$ on $\mathfrak{H}_{n}^{\mathbf{a}}$ with $X_{v},Y_{v}\in\mathrm{Her}_{n}(\mathbb{C})$ for each $v\in\mathbf{a}$. We identify $\mathrm{Her}_{n}(\mathbb{C})$ with $\mathbb{R}^{n^{2}}$ and define measures $dX_{v},dY_{v}$ as the standard measures on $\mathbb{R}^{n^{2}}$. We define a measure $dZ$ on $\mathfrak{H}_{n}^{\mathbf{a}}$ by
$dZ=\prod_{v\in\mathbf{a}}dX_{v}dY_{v}.$
For $F,G\in M_{\rho}(\widetilde{\Gamma}_{n},\omega)$, we define the Petersson inner product by
$(F,G)=\int_{D}\left\langle\rho(Y^{1/2},{}^{t}Y^{1/2})F(Z),\rho(Y^{1/2},{}^{t}Y^{1/2})G(Z)\right\rangle(\prod_{v\in\mathbf{a}}\det(Y_{v})^{-2n})dZ,$
where $Y=(Y_{v})_{v\in\mathbf{a}}=\mathrm{Im}(Z)$, $Y^{1/2}=(Y^{1/2}_{v})_{v\in\mathbf{a}}$ is a family of positive definite Hermitian matrices such that $(Y_{v}^{1/2})^{2}=Y_{v}$, and $D$ is a Siegel domain on $\mathfrak{H}_{n}^{\mathbf{a}}$ for $\widetilde{\Gamma}_{n}$. This integral converges if either $F$ or $G$ is a cusp form. We call a sequence of non-negative integers $\bm{k}=(k_{1},k_{2},\dots)$ a dominant integral weight if $k_{i}\geq k_{i+1}$ for all $i$ and $k_{i}=0$ for almost all $i$. The largest integer $m$ such that $k_{m}\neq 0$ is called the length of $\bm{k}$ and denoted by $\ell(\bm{k})$. Dominant integral weights of length at most $n$ correspond bijectively to the set of irreducible algebraic representations of $\mathrm{GL}_{n}(\mathbb{C})$. We denote by $(\rho_{n,\bm{k}},V_{n,\bm{k}})$ the irreducible representation of $\mathrm{GL}_{n}(\mathbb{C})$ corresponding to a dominant integral weight $\bm{k}$ with $\ell(\bm{k})\leq n$.
In what follows, we will frequently realize representation spaces in terms of bideterminants, so we briefly recall the relevant construction.
Let $\bm{k}=(k_{1},\dots,k_{n})$ be a dominant integral weight with $\ell(\bm{k})\leq n$. We realize the corresponding representation space $V_{n,\bm{k}}$ of $\mathrm{GL}_{n}(\mathbb{C})$ as a space of bideterminants (cf. Ibukiyama [13]).
Let $U=(u_{ij})$ be an $\ell(\bm{k})\times n$ matrix of variables. For a positive integer $a\leq\ell(\bm{k})$, let $SI_{n,a}$ denote the set of strictly increasing sequences $J=(j_{1},\dots,j_{a})$ of positive integers not exceeding $n$. For each $J\in SI_{n,a}$, define the submatrix $U_{J}$ of $U$ by
$U_{J}=\begin{pmatrix}u_{1,j_{1}}&\cdots&u_{1,j_{a}}\\ \vdots&\ddots&\vdots\\ u_{a,j_{1}}&\cdots&u_{a,j_{a}}\end{pmatrix}$.
A polynomial $P(U)$ in the entries of $U$ is called a bideterminant of weight $\bm{k}$ if it has the form
$P(U)=\prod_{i=1}^{m}\prod_{j=1}^{k_{i}-k_{i+1}}\det U_{J_{ij}},$
where $J_{ij}\in SI_{n,i}$ for each $j=1,\ldots,k_{i}-k_{i+1}$. Here we adopt the convention that the inner product is $1$ if $k_{i}=k_{i+1}$.
Let $BD_{\bm{k}}$ denote the set of all bideterminants of weight $\bm{k}$. For a commutative ring $R$ and an $R$-algebra $S$, let $S[U]_{\bm{k}}$ denote the $R$-module of all $S$-linear combinations of elements in $BD_{\bm{k}}$.
The group $\mathrm{GL}_{n}(\mathbb{C})$ acts on $\mathbb{C}[U]_{\bm{k}}$ via
$(g,P(U))\mapsto P(Ug),\quad g\in\mathrm{GL}_{n}(\mathbb{C}),\,P(U)\in\mathbb{C}[U]_{\bm{k}},$
and under this action, $\mathbb{C}[U]_{\bm{k}}$ provides a concrete realization of the representation space $V_{n,\bm{k}}$.
We define an inner product on $\mathbb{C}[U]_{\bm{k}}$ by
$\langle P(U),Q(U)\rangle=P(\partial_{U})\,\overline{Q}(U)\big|_{U=0},$
where $P(\partial_{U})$ denotes the differential operator obtained by replacing each variable $u_{ij}$ in $P(U)$ with $\frac{\partial}{\partial u_{ij}}$, and $\overline{Q}(U)$ is the polynomial obtained by taking the complex conjugates of the coefficients of $Q(U)$. Then we have
$\langle\rho_{n,\bm{k}}(g)P(U),Q(U)\rangle=\langle P(U),\rho_{n,\bm{k}}(g^{*})Q(U)\rangle$
for any $g\in\mathrm{GL}_{n}(\mathbb{C})$ and any $P(U),Q(U)\in\mathbb{C}[U]_{\bm{k}}$ (cf. [1]).
We fix a basis $\{v_{1},\ldots,v_{t}\}$ of the free module $\mathbb{Z}[U]_{\bm{k}}$; for example, one may take the basis associated with semi-standard Young tableaux (cf. [10, §4.5]). Let $K$ be a number field and let $\mathcal{O}$ denote the ring of integers of $K$. For a prime ideal $\mathfrak{p}$ of $\mathcal{O}$ and an element $a=\sum_{i=1}^{t}a_{i}v_{i}\in K[U]_{\bm{k}}$, we define the valuation of $a$ with respect to $\mathfrak{p}$ by
$v_{\mathfrak{p}}(a)=\min_{1\leq i\leq t}v_{\mathfrak{p}}(a_{i}).$
We say that $\mathfrak{p}$ divides $a$ if $v_{\mathfrak{p}}(a)>0$, and write $\mathfrak{p}|a$. This valuation is independent of the choice of the basis of $\mathbb{Z}[U]_{\bm{k}}$.
For a family $(\bm{k},\bm{l})=(\bm{k}_{v},\bm{l}_{v})_{v\in\mathbf{a}}$ of pairs of dominant integral weights such that $\ell(\bm{k}_{v})\leq n$ and $\ell(\bm{l}_{v})\leq n$ for any $v\in\mathbf{a}$, we define the representation $\rho_{n,(\bm{k},\bm{l})}=\operatorname*{\boxtimes}_{v\in\mathbf{a}}(\rho_{n,\bm{k}_{v}}\boxtimes\rho_{n,\bm{l}_{v}})$ of $\widetilde{\mathfrak{K}}_{n,\infty}^{\mathbb{C}}$. We put
$M_{(\bm{k},\bm{l})}(\widetilde{\Gamma}_{n},\omega)=M_{\rho_{n,(\bm{k},\bm{l})}}(\widetilde{\Gamma}_{n},\omega)\quad\text{and}\quad S_{(\bm{k},\bm{l})}(\widetilde{\Gamma}_{n},\omega)=S_{\rho_{n,(\bm{k},\bm{l})}}(\widetilde{\Gamma}_{n},\omega).$
When $\bm{k}=(\kappa_{v},\ldots,\kappa_{v})_{v\in\mathbf{a}}$ and $\bm{l}=(0,\ldots,0)_{v\in\mathbf{a}}$ for a family $\kappa=(\kappa_{v})_{v\in\mathbf{a}}$ of non-negative integers, we also put
${\det}^{\kappa}=\rho_{n,(\bm{k},\bm{l})},\quad M_{\kappa}(\widetilde{\Gamma}_{n},\omega)=M_{(\bm{k},\bm{l})}(\widetilde{\Gamma}_{n},\omega)\quad\text{and}\quad S_{\kappa}(\widetilde{\Gamma}_{n},\omega)=S_{(\bm{k},\bm{l})}(\widetilde{\Gamma}_{n},\omega).$
2.2. As functions on unitary groups over the adeles
Let $\mathfrak{K}_{n,\infty}$ be the stabilizer of $\bm{i}_{n}\in\mathfrak{H}_{n}^{\mathbf{a}}$ in $G_{n,\infty}$. Then $\mathfrak{K}_{n,\infty}$ is a maximal compact subgroup of $G_{n,\infty}$ (it is also a maximal compact subgroup of $\widetilde{G}^{+}_{n,\infty}$) and is isomorphic to $\prod_{v\in\mathbf{a}}\mathrm{U}(n)\times\mathrm{U}(n)$, which is given by
$\begin{array}[]{rccc}&\prod_{v\in\mathbf{a}}\mathrm{U}(n)\times\mathrm{U}(n)&\rightarrow&\mathfrak{K}_{n,\infty}\\ &(k_{1,v},k_{2,v})_{v\in\mathbf{a}}&\mapsto&\left(\mathfrak{c}\begin{pmatrix}k_{2,v}&0\\ 0&{}^{t}k_{1,v}^{-1}\end{pmatrix}\mathfrak{c}^{-1}\right)_{v\in\mathbf{a}},\end{array}$
where $\mathfrak{c}=\dfrac{1}{\sqrt{2}}\begin{pmatrix}1&\sqrt{-1}\\ \sqrt{-1}&1\\ \end{pmatrix}\in\mathrm{M}_{2n}(\mathbb{C})$.
The stabilizer of $\bm{i}_{n}$ in $\widetilde{G}^{+}_{n,\infty}$ is $\widetilde{\mathfrak{K}}^{+}_{n,\infty}=\prod_{v\in\mathbf{a}}(\mathbb{R}_{>0}\cdot\widetilde{\mathfrak{K}}_{n,\infty})$. Note that $\widetilde{\mathfrak{K}}^{+}_{n,\infty}$ is clearly not a maximal compact subgroup of $\widetilde{G}^{+}_{n,\infty}$.
For an infinite place $v\in\mathbf{a}$, we put $\mathfrak{g}_{n,v}=\mathrm{Lie}(\widetilde{G}_{n,v})$, $\mathfrak{k}_{n,v}=\mathrm{Lie}(\widetilde{\mathfrak{K}}_{n,v})$ and let $\mathfrak{g}^{\mathbb{C}}_{n,v}$ and $\mathfrak{k}^{\mathbb{C}}_{n,v}$ be the complexification of $\mathfrak{g}_{n,v}$ and $\mathfrak{k}_{n,v}$, respectively.
We have the Cartan decomposition $\mathfrak{g}_{n,v}=\mathfrak{k}_{n,v}\oplus\mathfrak{p}_{n,v}$.
We put
$\kappa_{v,ij}$ $=\mathfrak{c}\begin{pmatrix}0&0\\ 0&-e_{ji}\\ \end{pmatrix}\mathfrak{c}^{-1},\quad\kappa^{\prime}_{v,ij}=\mathfrak{c}\begin{pmatrix}e_{ij}&0\\ 0&0\\ \end{pmatrix}\mathfrak{c}^{-1},$
$\pi^{+}_{v,ij}$ $=\mathfrak{c}\begin{pmatrix}0&e_{ij}\\ 0&0\\ \end{pmatrix}\mathfrak{c}^{-1},\quad\mathrm{and}\quad\pi^{-}_{v,ij}=\mathfrak{c}\begin{pmatrix}0&0\\ e_{ij}&0\\ \end{pmatrix}\mathfrak{c}^{-1},$
where $e_{ij}\in\mathrm{M}_{n,n}(\mathbb{C})$ is the matrix whose only non-zero entry is 1 in $(i,j)$-component. $\{\kappa_{v,ij}\}$ is a basis of $\mathfrak{k}^{\mathbb{C}}_{n,v}$. Let $\mathfrak{p}^{+}_{n,v}$ (resp. $\mathfrak{p}^{-}_{n,v}$) be the $\mathbb{C}$-span of $\{\pi^{+}_{v,ij}\}$ (resp. $\{\pi^{-}_{v,ij}\}$) in $\mathfrak{g}^{\mathbb{C}}_{n,v}$. Then, we have $\mathfrak{p}^{\mathbb{C}}_{n,v}=\mathfrak{p}^{+}_{n,v}\oplus\mathfrak{p}^{-}_{n,v}\oplus\mathbb{C}\cdot I_{2n}$. We set $\mathfrak{g}_{n}=\prod_{v\in\mathbf{a}}\mathfrak{g}_{n,v}$, $\mathfrak{k}^{\mathbb{C}}_{n}=\prod_{v\in\mathbf{a}}\mathfrak{k}^{\mathbb{C}}_{n,v}$, etc.
Definition 2.2.
Let $\rho$ be a representation of $\widetilde{\mathfrak{K}}_{n,\infty}$ on a finite-dimensional complex vector space $V$, and let $\omega$ be a unitary character of $\mathbb{R}^{\mathbf{a}}$.
A Hermitian automorphic form on $\widetilde{G}^{+}_{n}(\mathbb{A})$ of weight $(\rho,V)$ and character $\omega$ is a $V$-valued smooth function $f:\widetilde{G}^{+}_{n}(\mathbb{A})\;\longrightarrow\;V$ satisfying the following conditions:
  • $f(ugk)=\omega(\nu(k_{\infty}))^{-1}\rho(\nu(k_{\infty})^{-1/2}k_{\infty})^{-1}f(g)$ for all $u\in\widetilde{G}^{+}_{n}$, $g\in\widetilde{G}^{+}_{n}(\mathbb{A})$, $k\in\widetilde{\mathfrak{K}}^{+}_{n,\mathbb{A}}$, where $k_{\infty}$ is the infinite part of $k$.
  • $f$ is of moderate growth.
  • $f$ is $Z(\mathfrak{g})$-finite, where $Z(\mathfrak{g}_{n})$ denotes the center of the universal enveloping algebra of $\mathfrak{g}_{n}$
We denote by $\widetilde{\mathcal{A}}_{n}(\rho,\omega)$ the complex vector space of Hermitian automorphic forms on $\widetilde{G}^{+}_{n}(\mathbb{A})$ of weight $\rho$ and character $\omega$.
Definition 2.3.
A Hermitian automorphic form $f\in\widetilde{\mathcal{A}}_{n}(\rho,\omega)$ is called a cusp form if
[…]
for any $g\in\widetilde{G}_{n}(\mathbb{A})$ and any unipotent radical $N$ of each proper parabolic subgroup of $\widetilde{G}_{n}$. We denote by $\widetilde{\mathcal{S}}_{n}(\rho,\omega)$ the complex vector space of cusp forms on $G_{n}(\mathbb{A})$ of weight $(\rho,V)$ and character $\omega$.
Remark 2.4.
When no character is specified, we simply write
$\widetilde{\mathcal{A}}_{n}(\rho)=\bigoplus_{\omega}\,\widetilde{\mathcal{A}}_{n}(\rho,\omega),\quad\text{and}\quad\widetilde{\mathcal{S}}_{n}(\rho)=\bigoplus_{\omega}\,\widetilde{\mathcal{S}}_{n}(\rho,\omega).$
We also define Hermitian automorphic forms for the unitary group $G_{n}$ in a similar way. Let $\mathcal{A}_{n}(\rho)$ (resp. $\mathcal{S}_{n}(\rho)$) denote the space of Hermitian automorphic forms (resp. cusp forms) on $G_{n}(\mathbb{A})$ of weight $(\rho,V)$.
For each finite place $v\in\mathbf{h}$, we take the Haar measure $dg_{v}$ on $G_{n,v}$, normalized so that the maximal compact subgroup $\mathfrak{K}_{n,v}$ has volume $1$.
For each archimedean place $v\in\mathbf{a}$, we choose the Haar measure $dg_{v}$ on $G_{n,v}$ so that the volume of $\mathfrak{K}_{n,v}$ is $1$, and the induced measure on the symmetric space $\mathfrak{H}_{n}\cong G_{n,v}/\mathfrak{K}_{n,v}$ is given by $(\det Y_{v})^{-2n}\,dZ_{v}$ on $\mathfrak{H}_{n}$.
By taking the restricted product, we obtain a Haar measure $dg=\prod_{v}dg_{v}$ on $G_{n}(\mathbb{A})$.
Using this choice, we define the Petersson inner product on the space of automorphic forms $\mathcal{A}_{n}(\rho)$ by
$(f,h)\;=\;\int_{G_{n}\backslash G_{n}(\mathbb{A})}\langle f(g),h(g)\rangle\,dg,$
for $f,h\in\mathcal{A}_{n}(\rho)$, where $dg$ denotes the Haar measure on $G_{n}\backslash G_{n}(\mathbb{A})$ induced from $dg$ on $G_{n}(\mathbb{A})$. If either $f$ or $h$ belongs to $\mathcal{S}_{n}(\rho)$, then the integral converges absolutely.
Let $U_{1}$ be the algebraic group defined by
$U_{1}(R)=\{\,x\in(K\otimes_{F}R)^{\times}\mid x\overline{x}=1\,\}$
for any $F$-algebra $R$, and let $\mathbb{G}_{m}$ denote the multiplicative group. By the strong approximation theorem for the special unitary group $\mathrm{SU}(n,n)$ and the short exact sequence
[…]
we obtain an isomorphism
[…]
$\cong$ $\;\bigl(F^{\times}\backslash\mathbb{A}^{\times}/\prod_{v\in\mathbf{h}}\mathcal{O}_{v}^{\times}\cdot\mathbb{A}^{\times}_{\infty}\bigr)\ltimes\bigl(K^{\times}\backslash\mathbb{A}_{K}^{\times}/\mathbb{A}^{\times}\prod_{v\in\mathbf{h}}\mathcal{O}_{K_{v}}^{\times}\mathbb{A}^{\times}_{K,\infty}\bigr),$ (2.3)
We now describe an explicit set of representatives for the relevant double coset space. Fix complete sets of representatives
$N_{0}:=\{\nu_{1}=1,\ldots,\nu_{h_{1}}\}\subset\mathbb{A}_{0}^{\times}\quad\text{for }F^{\times}\backslash\mathbb{A}^{\times}/\prod_{v\in\mathbf{h}}\mathcal{O}_{v}^{\times}\cdot\mathbb{A}^{\times}_{\infty},$
and
$Z_{0}:=\{z_{1}=1,\ldots,z_{h_{2}}\}\subset\mathbb{A}_{K,0}^{\times}\quad\text{for }K^{\times}\backslash\mathbb{A}_{K}^{\times}/\mathbb{A}^{\times}\prod_{v\in\mathbf{h}}\mathcal{O}_{K_{v}}^{\times}\mathbb{A}^{\times}_{K,\infty}.$
By the decomposition obtained above, each double coset in
[…]
is uniquely determined by a choice of parameters $\nu_{i}$ and $z_{j}$. Consequently, a complete set of double coset representatives may be indexed by pairs $(i,j)$. More precisely, we may choose a set of representatives $\gamma_{ij}$ in (2.4) as
$\mathfrak{G}_{0}=\mathfrak{G}^{(n)}_{0}:=\left\{\gamma_{ij}=\left(\mathbf{s}(\nu_{i})\,\mathbf{t}(z_{j})\right)\middle|\,\nu_{i}\in N_{0},\,z_{j}\in Z_{0}\right\}\subset\widetilde{G}^{+}_{n,0}$,
where
$\mathbf{s}(\nu)=\begin{pmatrix}\nu I_{n}&0\\ 0&I_{n}\end{pmatrix},\qquad\mathbf{t}(z)=\begin{pmatrix}I_{n-1}&0&0&0\\ 0&z&0&0\\ 0&0&I_{n-1}&0\\ 0&0&0&\overline{z}^{-1}\end{pmatrix}$.
We fix an ordering of the set of pairs $(i,j)$ and, with respect to this ordering, occasionally denote $\gamma_{ij}$ by $\gamma_{s}$.
In particular, the subset $\{\gamma_{1j}\}_{1\leq j\leq h_{2}}$ forms a complete set of representatives for the double coset
$G_{n}(F)\backslash G_{n}(\mathbb{A})/\mathfrak{K}_{n,\mathbb{A}}G_{n,\infty}.$
For each $i,j$, set
[…]
Define
$M_{\rho}(\widetilde{\Gamma}_{s},\omega)=\Bigl\{\,f:\mathfrak{H}_{n}^{\mathbf{a}}\to V\;\Bigm|\;f|_{(\rho,\omega)}\gamma=f\ \text{for all }\gamma\in\widetilde{\Gamma}_{s}\Bigr\},$
as before.
Given $(f_{1},\ldots,f_{h})\in\bigoplus_{i=1}^{h}M_{\rho}(\widetilde{\Gamma}_{i},\omega)$, we define a function
$(f_{1},\ldots,f_{h})^{\sharp}:\widetilde{G}^{+}_{n}(\mathbb{A})\longrightarrow V$
by
$(f_{1},\ldots,f_{h})^{\sharp}(g)=f_{s}\mid_{\rho,\omega}g_{\infty}(\bm{i}_{n})=\omega(\nu(g_{\infty}))^{-1}\,\rho(M(g_{\infty}))^{-1}\,f_{s}\bigl(g_{\infty}\langle\bm{i}_{n}\rangle\bigr),$ (2.5)
for
$g=u\gamma_{s}kg_{\infty},\quad u\in\widetilde{G}^{+}_{n}(F),\quad\gamma_{s}\in\mathfrak{G}_{0},\quad k\in\widetilde{\mathfrak{K}}^{+}_{n,0},\quad g_{\infty}\in\widetilde{G}^{+}_{n,\infty}$.
Then $(f_{1},\ldots,f_{h})^{\sharp}$ belongs to $\widetilde{\mathcal{A}}_{n}(\rho,\omega)$. We denote the image of $\bigoplus_{i=1}^{h}M_{\rho}(\widetilde{\Gamma}_{i},\omega)$ (resp. $\bigoplus_{i=1}^{h}S_{\rho}(\widetilde{\Gamma}_{i},\omega)$) under the map $\sharp$ in (2.5) by $\widetilde{\mathcal{A}}_{n}^{\sharp}(\rho,\omega)$ (resp. $\widetilde{\mathcal{S}}_{n}^{\sharp}(\rho,\omega)$). We also define $\mathcal{A}_{n}^{\sharp}(\rho)$ and $\mathcal{S}_{n}^{\sharp}(\rho)$ for $G_{n}$ in a similar way.
Let $\mathrm{Her}_{n}(K)$ be the set of all Hermitian matrices in $\mathrm{M}_{n}(K)$, and $\mathrm{Her}_{n}(\mathbb{A}_{K})$ be the set of all Hermitian matrices in $\mathrm{M}_{n}(\mathbb{A}_{K})$. We fix a non-trivial additive character $\mathbf{e}_{\mathbb{A}}(x)=\prod_{v\in\mathbf{h}}\psi_{v}(x_{v})\times\prod_{v\in\mathbf{a}}\mathbf{e}(x_{v})$ of $\mathbb{A}/F$, such that $\psi_{v}$ is trivial on $\mathcal{O}_{F_{v}}$ for any finite place $v\in\mathbf{h}$. For $S\in\mathrm{Her}_{n}(K)$, we define the $S$-th Fourier coefficient of $f\in\widetilde{\mathcal{A}}_{n}(\rho,\omega)$ by
[…]
for $g\in\widetilde{G}^{+}_{n}(\mathbb{A})$, where $\mathbf{n}(X)=\begin{pmatrix}I_{n}&X\\ 0&I_{n}\end{pmatrix}$. By the definition, we have the following lemma.
Lemma 2.5.
  1. (1)
    We have
    $f(g)=\sum_{S\in\mathrm{Her}_{n}(K)}A_{f}(g,S).$
  2. (2)
    For any $X\in\mathrm{Her}_{n}(\mathbb{A}\otimes_{F}K)$ and $k\in\widetilde{\mathfrak{K}}_{n,0}$ we have
    $A_{f}(\mathbf{n}(X)gk,S)=\mathbf{e}_{\mathbb{A}}(\operatorname{Tr}(SX))A_{f}(g,S).$
  3. (3)
    For any $\alpha\in\mathrm{GL}_{n}(K)$ and $\nu\in F_{>0}^{\times}$, we have
    $A_{f}(\mathbf{s}(\nu)\mathbf{m}(\alpha)g,S)=A_{f}(g,\nu\,S[\alpha]),$
    where $\mathbf{s}(\nu)=\begin{pmatrix}\nu I_{n}&0\\ 0&I_{n}\end{pmatrix}$, $\mathbf{m}(\alpha)=\begin{pmatrix}\alpha&0\\ 0&{\alpha^{*}}^{-1}\end{pmatrix}$, and $S[\alpha]={\alpha^{*}}S\alpha$.
We put
$\eta(g_{\infty},S)=\mathbf{e}(\operatorname{Tr}(Sg_{\infty}\left<\bm{i}_{n}\right>))\omega(\nu(g_{\infty}))^{-1}\rho(M(g_{\infty}))^{-1}$
for $S\in\mathrm{Her}_{n}(\mathbb{C})^{\mathbf{a}}$ and $g_{\infty}\in\widetilde{G}^{+}_{n,\infty}$.
Lemma 2.6.
For $g_{\infty}\in\widetilde{G}^{+}_{n,\infty}$, $X_{\infty}\in\mathrm{Her}_{n}(\mathbb{C})^{\mathbf{a}}$, $\nu_{\infty}\in(\mathbb{R}_{>0})^{\mathbf{a}}$ and $A_{\infty}\in\mathrm{GL}_{n}(\mathbb{C})^{\mathbf{a}}$, we have
$\eta(\mathbf{n}(X_{\infty})g_{\infty},S)$ $=\mathbf{e}(\operatorname{Tr}(SX_{\infty}))\eta(g_{\infty},S),$
$\eta(\mathbf{s}(\nu_{\infty})\mathbf{m}(A_{\infty})g_{\infty},S)$ $=\omega(\nu_{\infty})^{-1}\eta(g_{\infty},\nu_{\infty}S[A_{\infty}])\rho(\nu_{\infty}^{1/2}{A_{\infty}^{*}},\nu_{\infty}^{1/2}\,{}^{t}\!A_{\infty}).$
We note that for $g_{0}\in\widetilde{G}^{+}_{n,0}$, the Iwasawa decomposition gives
$g_{0}=\mathbf{n}(X_{0})\,\mathbf{s}(\nu_{0})\,\mathbf{m}(\alpha_{0})\,\mathbf{t}(x_{0})\,\gamma_{ij}\,k,$
for some $X_{0}\in\mathrm{Her}_{n}(\mathbb{A}_{0}\otimes_{F}K)$, $\nu\in F_{>0}^{\times}$, $\alpha\in\mathrm{GL}_{n}(K)$, $x_{0}\in\mathbb{A}$, $\gamma_{ij}\in\mathfrak{G}_{0}$, and $k\in\widetilde{\mathfrak{K}}_{n,0}$. Here $\nu_{0}$ and $\nu_{\infty}$ denote the finite and infinite components of $\nu\in F\subset\mathbb{A}$, respectively; the same convention applies to $\alpha$.
The two Fourier coefficients $a(f_{ij},T)$ and $A_{f}(g,S)$ for $f=(f_{1},\ldots,f_{h})^{\sharp}\in\widetilde{\mathcal{A}}_{n}^{\sharp}(\rho,\omega)$ introduced above are related as follows.
Proposition 2.7.
Let $g=g_{0}g_{\infty}\in\widetilde{G}^{+}_{n}(\mathbb{A})$ with $g_{0}=\mathbf{n}(X_{0})\,\mathbf{s}(\nu_{0})\,\mathbf{m}(\alpha_{0})\,\gamma_{ij}\,k\in\widetilde{G}^{+}_{n,0}$ as above (so $x_{0}=1$), and $g_{\infty}\in\widetilde{G}^{+}_{n,\infty}$. Then, for $f=(f_{11},\ldots,f_{h_{1}h_{2}})^{\sharp}\in\widetilde{\mathcal{A}}_{n}^{\sharp}(\rho,\omega)$, we have
$A_{f}(g,S)=\mathbf{e}_{\mathbb{A}}(\operatorname{Tr}(SX_{0}))\,\omega(\nu_{\infty})\,\eta(g_{\infty},S)\,\rho(\nu_{\infty}^{1/2}\,{\alpha_{\infty}^{*}},\nu_{\infty}^{1/2}\,{}^{t}\!\alpha_{\infty})^{-1}\,a(f_{ij},\nu S[\alpha]).$
In particular,
$a(f_{ij},S)=A_{f}(\gamma_{ij},S).$
Proof.
If we put
$f_{ij,\infty}(g_{\infty})=f(\gamma_{ij}g_{\infty})=f_{ij}|_{\rho,\omega}g_{\infty}(\bm{i}_{n}),$
for $g_{\infty}\in\widetilde{G}^{+}_{n,\infty}$, then we have
$f_{ij,\infty}(\mathbf{n}(X_{\infty})g_{\infty})=\sum_{T\in\mathrm{Her}_{n}(K)}\eta(g_{\infty},T)\,a(f_{ij},T)\,\mathbf{e}\!\left(\sum_{v\in\mathbf{a}}\operatorname{Tr}(T_{\sigma_{v}}X_{v})\right),$
for $X_{\infty}=(X_{v})_{v\in\mathbf{a}}\in\mathrm{Her}_{n}(\mathbb{C})^{\mathbf{a}}$ and $g_{\infty}\in\widetilde{G}^{+}_{n,\infty}$. On the other hand,
$f_{ij,\infty}(\mathbf{n}(X_{\infty})g_{\infty})$ $=\sum_{S\in\mathrm{Her}_{n}(K)}A_{f}(\gamma_{ij}\mathbf{n}(X_{\infty})g_{\infty},S)$
$=\sum_{S\in\mathrm{Her}_{n}(K)}A_{f}(\gamma_{ij}g_{\infty},S)\,\mathbf{e}\!\left(\sum_{v\in\mathbf{a}}\operatorname{Tr}(S_{\sigma_{v}}X_{v})\right).$
Comparing these two expressions, we obtain
$A_{f}(\gamma_{ij}g_{\infty},S)=\eta(g_{\infty},S)\,a(f_{ij},S),$
for $S\in\mathrm{Her}_{n}(K)$ and $g_{\infty}\in\widetilde{G}^{+}_{n,\infty}$.
If we put
$\mu(g)=\eta(g_{\infty},S)^{-1}A_{f}(g,S),$
then by Lemma 2.5 and Lemma 2.6, we have
$\mu(g)$ $=\mathbf{e}_{\mathbb{A}}(\operatorname{Tr}(SX_{0}))\,\eta(g_{\infty},S)^{-1}\,A_{f}(\mathbf{s}(\nu_{0})\mathbf{m}(\alpha_{0})\gamma_{ij}g_{\infty},S)$
$=\mathbf{e}_{\mathbb{A}}(\operatorname{Tr}(SX_{0}))\,\eta(g_{\infty},S)^{-1}\,A_{f}(\gamma_{ij}\mathbf{s}(\nu^{-1}_{\infty})\mathbf{m}(\alpha^{-1}_{\infty})g_{\infty},\nu S[\alpha])$
$=\mathbf{e}_{\mathbb{A}}(\operatorname{Tr}(SX_{0}))\,\eta(g_{\infty},S)^{-1}\,\eta(\mathbf{s}(\nu^{-1}_{\infty})\mathbf{m}(\alpha^{-1}_{\infty})g_{\infty},\nu_{\infty}S_{\infty}[\alpha_{\infty}])\,a(f_{ij},\nu S[\alpha])$
$=\mathbf{e}_{\mathbb{A}}(\operatorname{Tr}(SX_{0}))\,\omega(\nu_{\infty})\,\rho(\nu_{\infty}^{1/2}\,{\alpha_{\infty}^{*}},\nu_{\infty}^{1/2}\,{}^{t}\!\alpha_{\infty})^{-1}\,a(f_{ij},\nu S[\alpha]).$
Hence the assertion follows. ∎
3. Hecke operators on Hermitian automorphic forms
For a finite place $v\in\mathbf{h}$ corresponding to a prime ideal $\mathfrak{p}$ of $F$, let $\mathcal{H}_{n,\mathfrak{p}}$ be the convolution algebra of left and right $\widetilde{\mathfrak{K}}_{n,v}$-invariant compactly supported $\mathbb{Z}$-valued functions on $\widetilde{G}_{n,v}\cap\mathrm{M}_{2n}(\mathcal{O}_{K_{v}})$, which is called the spherical Hecke algebra at $\mathfrak{p}$. The elements of $\mathcal{H}_{n,\mathfrak{p}}$ are called Hecke operators at $\mathfrak{p}$. The spherical Hecke algebra $\mathcal{H}_{n,\mathfrak{p}}$ acts from the left on the space of automorphic forms $\widetilde{\mathcal{A}}_{n}(\rho_{n,(\bm{k},\bm{l})})$ as follows:
$(T\cdot f)(g)=\int_{\widetilde{G}_{n,v}\cap\mathrm{M}_{2n}(\mathcal{O}_{K_{v}})}f(gh)T(h^{-1})|\nu(h)|_{\mathfrak{p}}^{r_{n}}dh,$
where we put $r_{n}=r_{n,(\bm{k},\bm{l})}=\sum_{i=1}^{n}(k_{i}+l_{i})/2-n^{2}$, and $dh$ is a Haar measure on $\widetilde{G}_{n,v}\cap\mathrm{M}_{2n}(\mathcal{O}_{K_{v}})$ normalized so that the volume of $\widetilde{\mathfrak{K}}_{n,v}$ is 1.
Note that the characteristic functions $\mathbf{1}_{\widetilde{\mathfrak{K}}_{n,v}g\widetilde{\mathfrak{K}}_{n,v}}$, which generate the Hecke algebra, can be naturally identified with the corresponding double cosets $\widetilde{\mathfrak{K}}_{n,v}g\widetilde{\mathfrak{K}}_{n,v}$.
For $(f_{1},\ldots,f_{h})^{\sharp}$, the Hecke operator $T=\widetilde{\mathfrak{K}}_{n,v}g\widetilde{\mathfrak{K}}_{n,v}$ with $g\in\widetilde{G}_{n,v}$ acts as follows. If the double coset decomposes as
$\widetilde{\mathfrak{K}}_{n,v}g\widetilde{\mathfrak{K}}_{n,v}=\bigsqcup_{i\in I}\widetilde{\mathfrak{K}}_{n,v}g_{i},\quad g_{i}\in\widetilde{G}_{n,v},$
then there exist $u_{i,j}\in\widetilde{G}^{+}_{n}(F)$ and $s_{i,j}\in\{1,\ldots,h\}$ such that
$\gamma_{j}g_{i}\in u_{i,j}\gamma_{s_{i,j}}\widetilde{\mathfrak{K}}_{n,0}\widetilde{G}^{+}_{n,\infty}$
Accordingly, $T$ acts on $(f_{1},\ldots,f_{h})^{\sharp}$ as
$T\cdot(f_{1},\ldots,f_{h})^{\sharp}=\left(\sum_{i\in I}|N_{F/\mathbb{Q}}(u_{i,1}^{-1})|^{r_{n}}\cdot f_{s_{i,1}}|_{\rho,\omega}u_{i,1}^{-1},\;\ldots,\;\sum_{i\in I}|N_{F/\mathbb{Q}}(u_{i,h}^{-1})|^{r_{n}}\cdot f_{s_{i,h}}|_{\rho,\omega}u_{i,h}^{-1}\right)^{\sharp}.$
Definition 3.1.
We say that a continuous right $\widetilde{\mathfrak{K}}_{n,v}$-invariant function $f$ on $\widetilde{G}_{n,v}$ (or on $\widetilde{G}_{n}(\mathbb{A})$) is a $\mathfrak{p}$-Hecke eigenfunction if $f$ is an eigenfunction under the action of $\mathcal{H}_{n,\mathfrak{p}}$.
Definition 3.2.
We say that a Hermitian automorphic form $f\in\widetilde{\mathcal{A}}_{n}(\rho)$ (resp. $f\in\widetilde{\mathcal{S}}_{n}(\rho)$) is a Hecke eigenform (resp. a Hecke cusp form) if $f$ is a $\mathfrak{p}$-Hecke eigenfunction for any $\mathfrak{p}$.
The structure of the Hecke algebra has been investigated in detail by Raum [26].
Lemma 3.3 ([26, Lemma 3.1]).
Let $v$ be a finite place of $F$ corresponding to a prime ideal $\mathfrak{p}$ of $F$. Then the spherical Hecke algebra $\mathcal{H}_{n,\mathfrak{p}}=\mathcal{H}(\widetilde{G}_{n,v}\cap\mathrm{M}_{n}(\mathcal{O}_{K}),\,\widetilde{\mathfrak{K}}_{n,v})$ is described as follows:
  1. (1)
    If $\mathfrak{p}$ is inert in $K$, then $\mathcal{H}_{n,\mathfrak{p}}$ is generated by $\{\,T(\varpi),\ T_{1}(\varpi^{2}),\ \ldots,\ T_{n}(\varpi^{2})\,\}$, where
    $T(\varpi)=\widetilde{\mathfrak{K}}_{n,v}\;\mathrm{diag}(\varpi I_{n},I_{n})\;\widetilde{\mathfrak{K}}_{n,v},$
    and
    $T_{i}(\varpi^{2})=\widetilde{\mathfrak{K}}_{n,v}\;\mathrm{diag}(\varpi^{2}I_{n-i},\varpi I_{i},I_{n-i},\varpi I_{i})\;\widetilde{\mathfrak{K}}_{n,v}$
    for $1\leq i\leq n$. Here $\varpi$ denotes a uniformizer of $K_{\mathfrak{p}}$, and we assume $\varpi\in F$.
  2. (2)
    If $\mathfrak{p}$ is ramified in $K$ as $\mathfrak{P}^{2}$, then $\mathcal{H}_{n,\mathfrak{p}}$ is generated by $\{T_{0},T_{1},\ldots,T_{n}\}$, where
    $\qquad T_{i}(\varpi)=\widetilde{\mathfrak{K}}_{n,v}\;\mathrm{diag}(\varpi\overline{\varpi}I_{n-i},\varpi I_{i},I_{n-i},\varpi I_{i})\;\widetilde{\mathfrak{K}}_{n,v},$
    where $\varpi$ is a uniformizer of $K_{\mathfrak{P}}$ for $0\leq i\leq n$.
  3. (3)
    If $\mathfrak{p}$ splits in $K$ as $\mathfrak{P}\overline{\mathfrak{P}}$, then $\mathcal{H}_{n,\mathfrak{p}}\cong\mathcal{H}(\mathrm{Inv}_{n}^{\mathrm{Res}}(\mathcal{O}_{v}),\mathrm{GL}_{2n}(\mathcal{O}_{v}))$, where
    $\mathrm{Inv}_{n}^{\mathrm{Res}}(\mathcal{O}_{v})=\{(g,l)\in(\mathrm{GL}_{2n}(F_{v})\cap\mathrm{M}_{2n}(\mathcal{O}_{v}))\times\mathbb{Z}_{\geq 0}\mid\varpi^{l}g^{-1}\in\mathrm{M}_{2n}(\mathcal{O}_{v})\}$
    and $\mathrm{GL}_{2n}(\mathcal{O}_{v})$ acts trivially on the second component of $\mathrm{Inv}_{n}^{\mathrm{Res}}(\mathcal{O}_{v})$. In this case, $\mathcal{H}_{n,\mathfrak{p}}$ is generated by $\{T_{0}(\varpi),\ldots,T_{2n}(\varpi)\}$, where
    $T_{i}(\varpi)=\mathrm{GL}_{2n}(\mathcal{O}_{v})\;\mathrm{diag}(\varpi I_{i},I_{2n-i})\;\mathrm{GL}_{2n}(\mathcal{O}_{v})$
    for $0\leq i\leq 2n$. The second component of each $T_{i}(\varpi)$ is 1. Here, $\varpi$ is a uniformizer of $F_{v}$.
To study the action of these Hecke operators on $f=(f_{1},\ldots,f_{h})^{\sharp}\in\widetilde{\mathcal{A}}_{n}(\rho)$ more concretely, we now describe the left cosets of the generators introduced in the lemma above (cf. Freitag [9, IV, §3]).
For an inert or ramified place $v$ corresponding to a prime ideal $\mathfrak{p}$ of $F$, let $\mathfrak{m}\subset\widetilde{G}_{n,v}\cap\mathrm{M}_{n}(\mathcal{O}_{K_{v}})$ be a subset which is left and right $\widetilde{\mathfrak{K}}_{n,v}$-invariant and whose elements have the same similitude $\nu$. We set
$\mathfrak{A}(\mathfrak{m})=\left\{A\in\mathrm{GL}_{n}(K_{v})\cap\mathrm{M}_{n}(\mathcal{O}_{K_{v}})\middle|\begin{pmatrix}A&*\\ 0&\nu\cdot A^{*-1}\end{pmatrix}\in\mathfrak{m}\right\}.$
For $A\in\mathfrak{A}(\mathfrak{m})$, we define
$\mathfrak{B}(A,\mathfrak{m})=\left\{B\in\mathrm{M}_{n}(\mathcal{O}_{K_{v}})\middle|\begin{pmatrix}A&B\\ 0&\nu\cdot A^{*-1}\end{pmatrix}\in\mathfrak{m}\right\}.$
On $\mathfrak{B}(A,\mathfrak{m})$, we introduce an equivalence relation with respect to $D\in\mathrm{M}_{n}(\mathcal{O}_{K_{v}})$ by
$B\sim_{D}B^{\prime}\quad\Longleftrightarrow\quad B-B^{\prime}\in\mathrm{Her}_{n}(\mathcal{O}_{K_{v}})D.$
Lemma 3.4.
Assume that $v$ is inert or ramified in $K$, and let $\mathfrak{m}\subset\widetilde{G}_{n,v}\cap\mathrm{M}_{n}(\mathcal{O}_{K_{v}})$ be a subset which is left and right $\widetilde{\mathfrak{K}}_{n,v}$-invariant and whose elements have the same similitude $\nu$. Then a complete set of left coset representatives of $\widetilde{\mathfrak{K}}_{n,v}\backslash\mathfrak{m}$ is given by
[…]
In the rest of this section, we fix a uniformizer $\varpi$, and for each $i=1,\ldots,h_{1}$, choose a representative $\nu_{l_{i}}$ and an element $\pi_{i}\in F^{\times}$ such that
$\nu_{i}^{-1}\varpi\in\pi_{i}\nu_{l_{i}}^{-1}\left(\prod_{v\in\mathbf{h}}\mathcal{O}_{v}^{\times}\right)\mathbb{A}_{\infty}^{\times}.$
We fix $f=(f_{11},\ldots,f_{h_{1}h_{2}})^{\sharp}\in\widetilde{\mathcal{A}}_{n}(\rho,\omega)$ for the rest of this section.
3.1. The case where $v$ is inert
First, suppose that $v$ (and hence $\mathfrak{p}$) is inert in $K$. Define
$\tau^{(n)}_{s}(\varpi)=\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\,\mathrm{diag}(I_{n-s},\varpi I_{s})\,\mathrm{GL}_{n}(\mathcal{O}_{K_{v}}),$
and
$\tau^{(n)}_{s,t}(\varpi)=\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\,\mathrm{diag}(I_{n-s-t},\varpi I_{s},\varpi^{2}I_{t})\,\mathrm{GL}_{n}(\mathcal{O}_{K_{v}}).$
We put $\mathcal{O}_{K_{v}}^{(0)}=\{0\}$, and let $\mathcal{O}_{K_{v}}^{(l)}$ be a complete set of representatives for the residue classes of $\mathcal{O}_{K_{v}}/\varpi^{l}\mathcal{O}_{K_{v}}$ for any positive integer $l$. The following lemma follows by the same argument as in Freitag [9, IV, 2.7]. Let $q_{v}$ denote the cardinality of the residue field of $F_{v}$.
Lemma 3.5.
We have
$\mathfrak{A}\bigl(T(\varpi)\bigr)=\bigsqcup_{0\leq s\leq n}\tau^{(n)}_{s}(\varpi),\qquad\mathfrak{A}\bigl(T_{i}(\varpi^{2})\bigr)=\bigsqcup_{\begin{subarray}{c}s+t\leq n\\ s\geq i\end{subarray}}\tau^{(n)}_{s,t}(\varpi).$
Moreover, a complete set of representatives of $\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s}(\varpi)$ is given by
$\left\{A=\begin{pmatrix}\varpi^{k_{1}}&&a_{uv}\\ &\ddots&\\ &&\varpi^{k_{n}}\end{pmatrix}\;\middle|\;\begin{array}[]{l}k_{1}+\cdots+k_{n}=s,\quad k_{j}\in\{0,1\},\\[4.0pt] a_{uv}\in\mathcal{O}_{K_{v}}^{(k_{t})},\;1\leq u<v\leq n,\\[4.0pt] a_{uv}=0\text{ if }k_{u}=k_{v}=1\end{array}\right\}.$
Remark 3.6.
If we fix $k_{1},\ldots,k_{n}\in\{0,1\}$, then the number of matrices $A$ satisfying the above conditions is $q_{v}^{-s(s+1)}\;\prod_{j=1}^{n}q_{v}^{2jk_{j}}$.
Using Lemma 3.4, we obtain the following counting result.
Lemma 3.7.
  1. (1)
    A complete set of representatives of $\mathfrak{B}(\mathrm{diag}(I_{n-s},\varpi I_{s}),T(\varpi))/\!\sim_{D}$ is given by
    $\left\{B=\begin{pmatrix}B_{0}&0\\ 0&0\end{pmatrix}\,\middle|\,B_{0}\in\mathrm{Her}_{n-s}(\mathcal{O}_{K_{v}})/\varpi\mathrm{Her}_{n-s}(\mathcal{O}_{K_{v}})\right\}.$
    In particular, for any $A\in\tau^{(n)}_{s}(\varpi)$, we have
    $\bigl(\mathfrak{B}(A,T(\varpi))/\!\sim_{D}\bigr)=q_{v}^{(n-s)^{2}}.$
  2. (2)
    A complete set of representatives of $\mathfrak{B}(\mathrm{diag}(I_{n-s-t},\varpi I_{s},\varpi^{2}I_{t}),T_{i}(\varpi^{2}))/\!\sim_{D}$ is given by
    $\left\{B=\begin{pmatrix}B_{1}&B_{2}&0\\[6.0pt] \varpi{B_{2}^{*}}&\ \begin{matrix}B_{3}\ &0\\[6.0pt] 0\ &0\end{matrix}\ &0\\[6.0pt] 0&0&0\end{pmatrix}\,\middle|\,\begin{array}[]{l}B_{1}\in\mathrm{Her}_{n-s-t}(\mathcal{O}_{K_{v}})/\varpi^{2}\mathrm{Her}_{n-s-t}(\mathcal{O}_{K_{v}}),\\ B_{2}\in\mathrm{M}_{n-s-t,s}(\mathcal{O}_{K_{v}})/\varpi\mathrm{M}_{n-s-t,s}(\mathcal{O}_{K_{v}}),\\ B_{3}\in\mathrm{Her}_{s-i}(\mathcal{O}_{K_{v}})/\varpi\mathrm{Her}_{s-i}(\mathcal{O}_{K_{v}}),\\ \det(B_{3})\not\in(\varpi)\end{array}\right\}.$
By combining the above lemmas, we obtain the following double coset decomposition:
$T(\varpi)=\bigsqcup_{\begin{subarray}{c}0\leq s\leq n\\[2.0pt] A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s}(\varpi)\\[2.0pt] B\in\mathfrak{B}(A,T(\varpi))/\!\sim_{D}\end{subarray}}\widetilde{\mathfrak{K}}_{n,v}\begin{pmatrix}A&B\\ 0&\varpi(A^{*})^{-1}\end{pmatrix}$.
Hence, for any $f\in\widetilde{\mathcal{A}}_{n}(\rho_{n,(\bm{k},\bm{l})})$ we have
$(T\cdot f)(g)=q_{v}^{r_{n}}\,\sum_{s,A,B}f\!\left(g\begin{pmatrix}A&B\\ 0&\varpi(A^{*})^{-1}\end{pmatrix}^{-1}\right).$
For each $\nu_{i}\in N_{0}$, choose a representative $\nu_{l_{i}}\in N_{0}$ and an element $\pi_{i}\in F^{\times}$ such that
$\nu_{i}^{-1}\varpi\in\pi_{i}\nu_{l_{i}}^{-1}\left(\prod_{v\in\mathbf{h}}\mathcal{O}_{v}^{\times}\right)\mathbb{A}_{\infty}^{\times}.$
For each $A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s}(\varpi)$, there exists $\alpha_{A}\in\mathrm{GL}_{n}(K)$ such that
$\varpi\,\mathbf{t}_{0}(z_{j})\,A^{-1}\in\alpha_{A}\,\mathbf{t}_{0}\!\bigl(z_{j}(\nu_{i}\nu_{l_{i}}^{-1})^{\,n-s}\bigr)\,\Bigl(\prod_{v\in\mathbf{h}}\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\Bigr)\,\mathrm{GL}_{n}(\mathbb{A}_{K,\infty})$
and $\det\alpha_{A}=\pi_{i}^{n-s}$, where we set
$\mathbf{t}_{0}(z)=\begin{pmatrix}I_{n-1}&0\\[2.0pt] 0&z\end{pmatrix}\quad(z\in K),$
so that $\mathbf{t}(z)=\mathbf{m}(\mathbf{t}_{0}(z))$.
Under this choice, we obtain
$\gamma_{ij}\begin{pmatrix}A&B\\ 0&\varpi{A^{*}}^{-1}\end{pmatrix}^{-1}$ $=\gamma_{ij}\mathbf{n}(-A^{-1}B)\mathbf{s}(\varpi^{-1})\mathbf{m}(\varpi A^{-1})$
$=\mathbf{n}\!\left(-\nu_{i}(A^{-1}B)[\mathbf{t}_{0}(\overline{z_{j}})]\right)\mathbf{s}(\nu_{i}\varpi^{-1})\mathbf{t}(z_{j})\mathbf{m}(\varpi A^{-1})$
$\in\mathbf{n}\!\left(-\nu_{i}(A^{-1}B)[\mathbf{t}_{0}(\overline{z_{j}})]\right)\mathbf{s}(\pi_{i,0}^{-1})\mathbf{m}(\alpha_{A,0})\gamma_{l_{i}j}\mathbf{t}\!\bigl((\nu_{i}\nu_{l_{i}}^{-1})^{n-s}\bigr)\widetilde{\mathfrak{K}}_{n,0},$
where $\pi_{i,0}$ and $\alpha_{A,0}$ denote the finite components of $\pi_{i}$ and $\alpha_{A}$, respectively.
Assume that
$\mathbf{t}\!\bigl((\nu_{i}\nu_{l_{i}}^{-1})^{n-s}\bigr)\in x_{is}\,\widetilde{\mathfrak{K}}_{n,0}\,h_{is},$
with $x_{is}\in\widetilde{G}_{n}^{+}$ and $h_{is}\in\widetilde{G}_{n,\infty}^{+}$. Then, if we put $q_{\pi_{i}}=N_{F/\mathbb{Q}}(\pi_{i})$, we have
$A_{T(\varpi)\cdot f}(\gamma_{ij},S)$ $=q_{\pi_{i}}^{r_{n}}\sum_{s=0}^{n}\sum_{A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s}(\varpi)}\left(\sum_{B}\mathbf{e}_{\mathbb{A}}\!\left(-\operatorname{Tr}\!\left(S\nu_{i}(A^{-1}B)[\mathbf{t}_{0}(\overline{z_{j}})]\right)\right)\right)$
$\quad\cdot\omega(\pi_{i,\infty}^{-1})\rho(\pi_{i,\infty}^{1/2}\alpha^{*-1}_{A,\infty},\pi_{i,\infty}^{1/2}\,{}^{t}\!\alpha^{-1}_{A,\infty})a\!\left(f_{l_{i}j}|_{\rho,\omega}h_{is},\pi_{i}^{-1}S[\alpha_{A}]\right)$
$=q_{\pi_{i}}^{2r_{n}+n^{2}}\sum_{s=0}^{n}\sum_{A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s}(\varpi)}\left(\sum_{B}\mathbf{e}_{\mathbb{A}}\!\left(-\operatorname{Tr}\!\left(\nu_{i}S[\mathbf{t}_{0}(z_{j})](A^{-1}B)\right)\right)\right)$
$\quad\cdot\omega(\pi_{i,\infty}^{-1})\rho(\alpha^{*-1}_{A,\infty},\,^{t}\!\alpha^{-1}_{A,\infty})a\!\left(f_{l_{i}j}|_{\rho,\omega}h_{is},\pi_{i}^{-1}S[\alpha_{A}]\right)$
by Proposition 2.7. Define
$Q_{s}(S)=\left\{A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s}(\varpi)\;\middle|\;\sum_{B\in\mathfrak{B}(A,T(\varpi))/\!\sim_{D}}\mathbf{e}_{\mathbb{A}}\!\big(-\operatorname{Tr}(SA^{-1}B)\big)\neq 0\right\}.$
By Lemma 3.7, we obtain the following formula:
$A_{T(\varpi)\cdot f}(\gamma_{ij},S)=q_{\pi_{i}}^{2r_{n}+n^{2}}\sum_{s=0}^{n}q_{v}^{(n-s)^{2}}\sum_{A\in Q_{s}(\nu_{i}S[\mathbf{t}_{0}(z_{j})])}$ $\omega(\pi_{i,\infty}^{-1})\,\rho(\alpha^{*-1}_{A,\infty},\,^{t}\!\alpha^{-1}_{A,\infty})\,$
$\quad\cdot a\!\left(f_{i1}|_{\rho,\omega}h_{is},\pi_{i}^{-1}S[\alpha_{A}]\right).$ (3.1)
The operator $T_{i}(\varpi^{2})$ can be treated in a similar manner. For each $\nu_{i}\in N_{0}$, choose a representative $\nu_{l^{\prime}_{i}}\in N_{0}$ and an element $\pi_{i}^{(2)}\in F^{\times}$ such that
$\nu_{i}^{-1}\varpi^{2}\in\pi_{i}^{(2)}\nu_{l^{\prime}_{i}}^{-1}\left(\prod_{v\in\mathbf{h}}\mathcal{O}_{v}^{\times}\right)\mathbb{A}_{\infty}^{\times}.$
Note that the subscript $(2)$ is only notational and does not indicate squaring. For each $A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s,t}(\varpi)$, there exists $\alpha_{A}\in\mathrm{GL}_{n}(K)$ such that
$\varpi^{2}\mathbf{t}_{0}(z_{j})A^{-1}\in\alpha_{A}\,\mathbf{t}_{0}\!\bigl(z_{j}(\nu_{i}\nu_{l_{i}}^{-1})^{2n-2t-s}\bigr)\Bigl(\prod_{v\in\mathbf{h}}\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\Bigr)\mathrm{GL}_{n}(\mathbb{A}_{K,\infty})$
and $\det\alpha_{A}=\pi_{i}^{2n-2t-s}$. Assume that
$\mathbf{t}\!\bigl((\nu_{i}\nu_{l_{i}}^{-1})^{2n-2t-s}\bigr)\in y_{is}\,\widetilde{\mathfrak{K}}_{n,0}h^{\prime}_{is},$
with $y_{is}\in\widetilde{G}_{n}^{+}$ and $h^{\prime}_{is}\in\widetilde{G}_{n,\infty}^{+}$. Then, if we put $q_{\pi_{i}^{(2)}}=N_{F/\mathbb{Q}}(\pi_{i}^{(2)})$, we obtain
$A_{T_{i}(\varpi^{2})\cdot f}(\gamma_{ij},S)$ $=q_{\pi_{i}^{(2)}}^{2r_{n}+n^{2}}\sum_{\begin{subarray}{c}s+t\leq n\\ s\geq i\end{subarray}}\sum_{A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s,t}(\varpi)}\left(\sum_{B\in\mathfrak{B}(A,T(\varpi))/\!\sim_{D}}\mathbf{e}_{\mathbb{A}}\!\left(-\operatorname{Tr}\!\big(\nu_{i}S[\mathbf{t}_{0}(z_{j})](A^{-1}B)\big)\right)\right)$
$\quad\cdot\omega((\pi^{(2)}_{i,\infty})^{-1})\rho(\alpha^{*-1}_{A,\infty},\,{}^{t}\!\alpha^{-1}_{A,\infty})a\!\left(f_{l^{\prime}_{i}j}|_{\rho,\omega}h^{\prime}_{is},(\pi^{(2)}_{i,\infty})^{-1}S[\alpha_{A}]\right).$
The summation over $B$ is generally difficult to evaluate directly, unlike in the case of $T(\varpi)$. Therefore, following [11] and [4], we introduce a new operator $\widetilde{T_{i}}(\varpi^{2})$ as a linear combination of several $T_{i}(\varpi^{2})$.
Definition 3.8.
For $1\leq i\leq n$, we define the Hecke operator
$\widetilde{T_{i}}(\varpi^{2})=\sum_{j=1}^{i}\begin{bmatrix}n-j\\ i-j\end{bmatrix}_{q_{v}^{2}}T_{j}(\varpi^{2}),$
where
$\begin{bmatrix}s\\ t\end{bmatrix}_{q}=\prod_{k=1}^{t}\frac{q^{s-k+1}-1}{q^{k}-1}$
denotes the $q$-binomial coefficient.
By the same argument as in Proposition 5.1 of [4], we then obtain
$A_{\widetilde{T_{i}}(\varpi^{2})\cdot f}(\gamma_{ij},S)=$ $q_{\pi_{i}^{(2)}}^{2r_{n}+n^{2}}\sum_{\begin{subarray}{c}s+t\leq n\\ s\geq i\end{subarray}}\sum_{A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s,t}(\varpi)}\left(\sum_{B}\mathbf{e}_{\mathbb{A}}\!\left(-\operatorname{Tr}\!\big(\nu_{i}S[\mathbf{t}_{0}(z_{j})](A^{-1}B)\big)\right)\right)$
$\cdot\omega((\pi^{(2)}_{i,\infty})^{-1})\rho(\alpha^{*-1}_{A,\infty},\,^{t}\!\alpha_{A,\infty}^{-1})a\!\left(f_{l^{\prime}_{i}j}|_{\rho,\omega}h^{\prime}_{is},(\pi^{(2)}_{i,\infty})^{-1}S[\alpha_{A}]\right),$
where $B$ runs through a complete set of representatives of
$\mathfrak{B}(A,s,t):=\left\{AH\;\middle|\;H=\begin{pmatrix}H_{1}&H_{2}&0\\[6.0pt] \varpi{H_{2}^{*}}&H_{3}&0\\[6.0pt] 0&0&0\end{pmatrix},\quad\begin{array}[]{l}H_{1}\in\mathrm{Her}_{n-s-t}(\mathcal{O}_{K_{v}})/\varpi^{2}\mathrm{Her}_{n-s-t}(\mathcal{O}_{K_{v}}),\\[4.0pt] H_{2}\in\mathrm{M}_{n-s-t,s}(\mathcal{O}_{K_{v}})/\varpi\mathrm{M}_{n-s-t,s}(\mathcal{O}_{K_{v}}),\\[4.0pt] H_{3}\in\mathrm{Her}_{s}(\mathcal{O}_{K_{v}})/\varpi\mathrm{Her}_{s}(\mathcal{O}_{K_{v}})\end{array}\right\}.$
If we define
$Q_{st}(S)=\left\{A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s,t}(\varpi)\;\middle|\;\sum_{B\in\mathfrak{B}(A,s,t)}\mathbf{e}_{\mathbb{A}}\!\big(-\operatorname{Tr}(SA^{-1}B)\big)\neq 0\right\},$
then it follows that
$A_{\widetilde{T_{i}}(\varpi^{2})\cdot f}(\gamma_{ij},S)$ $=q_{\pi_{i}^{(2)}}^{2r_{n}+n^{2}}\sum_{\begin{subarray}{c}s+t\leq n\\ s\geq i\end{subarray}}q_{v}^{2(n-s-t)(n-t)+s^{2}}\sum_{A\in Q_{st}(\nu_{i}S[\mathbf{t}_{0}(z_{j})])}\omega((\pi^{(2)}_{i,\infty})^{-1})$
$\quad\cdot\rho(\alpha^{*-1}_{A,\infty},\,{}^{t}\!\alpha_{A,\infty}^{-1})a\!\left(f_{l^{\prime}_{i}j}|_{\rho,\omega}h^{\prime}_{is},(\pi^{(2)}_{i,\infty})^{-1}S[\alpha_{A}]\right),$ (3.2)
since $\mathfrak{B}(A,s,t)$ contains $q_{v}^{2(n-s-t)(n-t)+s^{2}}$ elements.
3.2. The case where $v$ is ramified
Next, we consider the case where $v$ (and hence $\mathfrak{p}$) is ramified in $K$. This case can be treated in a way completely parallel to the inert case, but we include full details for completeness.
Define
$\tau^{(n)}_{s,t}(\varpi)=\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\,\mathrm{diag}(I_{n-s-t},\,\varpi I_{s},\,\varpi\overline{\varpi}I_{t})\,\mathrm{GL}_{n}(\mathcal{O}_{K_{v}}).$
Choose representative $\nu_{l_{i}}\in N_{0}$ and an element $\pi_{i}\in F^{\times}$ such that
$\nu_{i}^{-1}\varpi\overline{\varpi}\in\pi_{i}\nu_{l_{i}}^{-1}\left(\prod_{v\in\mathbf{h}}\mathcal{O}_{v}^{\times}\right)\mathbb{A}_{\infty}^{\times}.$
For each $A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s,t}(\varpi)$ and representative $z_{j}\in Z_{0}$, there exists representative $z_{m_{sj}}\in Z_{0}$, $\nu_{stj}\in N_{0}$ and $\alpha_{A}\in\mathrm{GL}_{n}(K)$ such that
$\varpi\overline{\varpi}\,\mathbf{t}_{0}(z_{j})\,A^{-1}\in\alpha_{A}\,\mathbf{t}_{0}\bigl(z_{sj}\nu_{stj}\bigr)\,\Bigl(\prod_{v\in\mathbf{h}}\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\Bigr)\,\mathrm{GL}_{n}(\mathbb{A}_{K,\infty}).$
Assume that
$\mathbf{t}\!\bigl(\nu_{stj}\bigr)\in y_{stj}\,\widetilde{\mathfrak{K}}_{n,0}h^{\prime}_{stj},$
with $y_{stj}\in\widetilde{G}_{n}^{+}$ and $h^{\prime}_{stj}\in\widetilde{G}_{n,\infty}^{+}$.
Definition 3.9.
For $0\leq i\leq n$, define the Hecke operator
$\widetilde{T_{i}}(\varpi)=\sum_{j=0}^{i}\begin{bmatrix}n-j\\ i-j\end{bmatrix}_{q_{v}}T_{j}(\varpi).$
Now define the auxiliary set
$\mathfrak{B}(A,s,t):=\left\{AH\;\middle|\;H=\begin{pmatrix}H_{1}&H_{2}&0\\[6.0pt] \overline{\varpi}{H_{2}^{*}}&H_{3}&0\\[6.0pt] 0&0&0\end{pmatrix},\;\begin{array}[]{l}H_{1}\in\mathrm{Her}_{n-s-t}(\mathcal{O}_{K_{v}})\big/\varpi\overline{\varpi}\mathrm{Her}_{n-s-t}(\mathcal{O}_{K_{v}}),\\[4.0pt] H_{2}\in\mathrm{M}_{n-s-t,s}(\mathcal{O}_{K_{v}})\big/\varpi\mathrm{M}_{n-s-t,s}(\mathcal{O}_{K_{v}}),\\[4.0pt] H_{3}\in\mathrm{Her}^{-}_{s}(\mathcal{O}_{K_{v}})\big/\varpi\mathrm{Her}_{s}(\mathcal{O}_{K_{v}})\end{array}\right\},$
where $\mathrm{Her}^{-}_{s}(\mathcal{O}_{K_{v}})$ denotes the $\mathcal{O}_{v}$-module of anti-Hermitian matrices of size $s$, that is,
$\mathrm{Her}^{-}_{s}(\mathcal{O}_{K_{v}})=\{\,X\in\mathrm{M}_{s}(\mathcal{O}_{K_{v}})\mid X=-{X^{*}}\,\}.$
Define
$Q_{st}(S)=\left\{A\in\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\backslash\tau^{(n)}_{s,t}(\varpi)\;\middle|\;\sum_{B\in\mathfrak{B}(A,s,t)}\mathbf{e}_{\mathbb{A}}\!\bigl(-\operatorname{Tr}(SA^{-1}B)\bigr)\neq 0\right\}.$
Then, if we put $q_{\pi_{i}}=N_{F/\mathbb{Q}}(\pi_{i})$, we have
$A_{\widetilde{T_{i}}(\varpi)\cdot f}(\gamma_{ij},S)$ $=q_{\pi_{i}}^{2r_{n}+n^{2}}\sum_{\begin{subarray}{c}s+t\leq n\\ s\geq i\end{subarray}}q_{v}^{(n-s-t)(n-t)+s(s-1)/2}\sum_{A\in Q_{st}(\nu_{i}S[\mathbf{t}_{0}(z_{j})])}\omega(\pi_{i,\infty}^{-1})$
$\qquad\cdot\rho(\alpha^{*-1}_{A,\infty},\,^{t}\!\alpha_{A,\infty}^{-1})a\!\left(f_{l^{\prime}_{i}m_{sj}}|_{\rho,\omega}h^{\prime}_{stj},\pi_{i}^{-1}S[\alpha_{A}]\right).$ (3.3)
3.3. The case where $v$ is split
By the same argument as in Lemma 3.5, we have the following lemma.
Lemma 3.10.
A complete set of representatives of $\mathrm{GL}_{2n}(\mathcal{O}_{v})\backslash T_{i}(\varpi)$ is given by
$\mathfrak{S}_{i}(\varpi)=\left\{\begin{pmatrix}\varpi^{k_{1}}&&a_{uv}\\ &\ddots&\\ &&\varpi^{k_{2n}}\end{pmatrix}\;\middle|\;\begin{array}[]{l}k_{1}+\cdots+k_{2n}=i,\quad k_{j}\in\{0,1\},\\[4.0pt] a_{uv}\in\mathcal{O}^{(k_{v})}_{v},\;1\leq u<v\leq 2n,\\[4.0pt] a_{uv}=0\ \text{if}\ k_{u}=k_{v}=1\end{array}\right\},$
where $\mathcal{O}^{(0)}_{v}=\{0\}$ and $\mathcal{O}^{(1)}_{v}\subset\mathcal{O}_{v}$ is a fixed set of representatives of $\mathcal{O}_{v}/\varpi\mathcal{O}_{v}$.
For each $\nu_{i}\in N_{0}$, choose a representative $\nu_{l_{i}}\in N_{0}$ and an element $\pi_{i}\in F^{\times}$ such that
$\nu_{i}^{-1}\varpi\in\pi_{i}\,\nu_{l_{i}}^{-1}\Bigl(\prod_{v\in\mathbf{h}}\mathcal{O}_{v}^{\times}\Bigr)\mathbb{A}_{\infty}^{\times}.$
Put
$\mathfrak{A}_{s}(\varpi):=\left\{A=\begin{pmatrix}\varpi^{k_{1}}&&a_{uv}\\ &\ddots&\\ &&\varpi^{k_{n}}\end{pmatrix}\;\middle|\;k_{1}+\cdots+k_{n}=s\right\}.$
We view $\mathfrak{A}_{s}(\varpi)$ as the set of upper-left $n\times n$ blocks of matrices in $\mathfrak{S}_{i}(\varpi)$.
Then for each $A\in\mathfrak{A}_{s}(\varpi)$, $D\in\mathfrak{A}_{i-s}(\varpi)$ and representative $z_{j}\in Z_{0}$, there exist $z_{m_{ij}}\in Z_{0}$ and $\alpha_{A}\in\mathrm{GL}_{n}(K)$ such that
$\mathbf{t}_{0}(z_{j})\,(\varpi A^{-1},\,{}^{t}D)\in\alpha_{A}\,\mathbf{t}_{0}(z_{m_{ij}}(\nu_{i}\nu_{l_{i}}^{-1})^{i-s})\,\Bigl(\prod_{v\in\mathbf{h}}\mathrm{GL}_{n}(\mathcal{O}_{K_{v}})\Bigr)\,\mathrm{GL}_{n}(\mathbb{A}_{K,\infty}).$
Here $(\varpi A^{-1},\,{}^{t}D)$ denotes the element of $\mathrm{GL}_{n}(\mathbb{A}_{K})$ whose $\mathfrak{P}$-component is $\varpi A^{-1}$, whose $\overline{\mathfrak{P}}$-component is $^{t}D$, and whose other components are the identity matrix.
Assume that
$\mathbf{t}\!\bigl((\nu_{i}\nu_{l_{i}}^{-1})^{i-s}\bigr)\in y_{is}\,\widetilde{\mathfrak{K}}_{n,0}h^{\prime}_{is},$
with $y_{is}\in\widetilde{G}_{n}^{+}$ and $h^{\prime}_{is}\in\widetilde{G}_{n,\infty}^{+}$.
Define
$\mathfrak{B}(A,D)=\left\{B\;\middle|\;\begin{pmatrix}A&B\\ 0&D\end{pmatrix}\in\mathfrak{S}_{i}(\varpi)\right\}$
for each $A\in\mathfrak{A}_{s}(\varpi)$ and $D\in\mathfrak{A}_{i-s}(\varpi)$ and put
$Q_{s}(S)=\left\{(A,D)\in\mathfrak{A}_{s}(\varpi)\times\mathfrak{A}_{i-s}(\varpi)\;\middle|\;\sum_{B\in\mathfrak{B}(A,D)}\mathbf{e}_{\mathbb{A}}\!\bigl(-\operatorname{Tr}(SA^{-1}B)\bigr)\neq 0\right\}.$
Then, if we put $q_{\pi_{i}}=N_{F/\mathbb{Q}}(\pi_{i})$, we have
$A_{T_{i}(\varpi)\cdot f}(\gamma_{ij},S)$ $=q_{\pi_{i}}^{2r_{n}+n^{2}}\sum_{s=0}^{i}\sum_{(A,D)\in Q_{s}(\nu_{i}S[\mathbf{t}_{0}(z_{j})])}n(A,D)\omega(\pi_{i,\infty}^{-1})$
$\quad\cdot\rho(\alpha^{*-1}_{A,\infty},\,{}^{t}\!\alpha_{A,\infty}^{-1})a\!\left(f_{l_{i},m_{ij}}|_{\rho,\omega}h^{\prime}_{is},\pi_{i}^{-1}S[\alpha_{A}]\right),$ (3.4)
where we put
$n(A,D)=q_{v}^{-i(i+1)/2}\;\prod_{j=1}^{2n}q_{v}^{jk_{j}}-q_{v}^{-s(s+1)/2}\;\prod_{j=1}^{n}q_{v}^{jk_{j}}-q_{v}^{-(i-s)(i-s+1)/2}\;\prod_{j=n+1}^{2n}q_{v}^{jk_{j}}$
for
$A=\begin{pmatrix}\varpi^{k_{1}}&&a_{uv}\\ &\ddots&\\ &&\varpi^{k_{n}}\end{pmatrix},\quad D=\begin{pmatrix}\varpi^{k_{n+1}}&&a_{uv}\\ &\ddots&\\ &&\varpi^{k_{2n}}\end{pmatrix}$.
4. Pullback formula
4.1. Differential operators on Hermitian automorphic forms
Let $n_{1},n_{2}$ be positive integers such that $n_{1}\geq n_{2}\geq 1$ and put $n=n_{1}+n_{2}$. We diagonally embed $\mathfrak{H}_{n_{1}}^{\mathbf{a}}\times\mathfrak{H}_{n_{2}}^{\mathbf{a}}\hookrightarrow\mathfrak{H}_{n}^{\mathbf{a}}$ and define the diagonal embedding $\iota_{n_{1},n_{2}}:G_{n_{1}}\times G_{n_{2}}\hookrightarrow G_{n}$.
Let $(\rho_{s}\boxtimes\tau_{s},V_{s})$ be a representation of $\mathfrak{K}_{n_{s},\infty}^{\mathbb{C}}$ for $s=1,2$. We fix tuples $\kappa=(\kappa_{v})_{v\in\mathbf{a}},\,\nu=(\nu_{v})_{v\in\mathbf{a}}$ of positive integers.
We will consider $V:=V_{1}\otimes V_{2}$-valued differential operators $\mathbb{D}$ on $\mathcal{A}_{n}^{\sharp}({\det}^{\kappa})$, satisfying Condition (A) below:
Condition (A).
For any automorphic form $f\in\mathcal{A}_{n}^{\sharp}(\det\nolimits^{\kappa}\boxtimes\det^{\nu})$ , we have
$\mathbb{D}f(\iota_{n_{1},n_{2}}(g_{1},g_{2}))\in\mathcal{A}_{n_{1}}^{\sharp}(\det\nolimits^{\kappa}\rho_{1}\boxtimes\det\nolimits^{\nu}\tau_{1})\otimes\mathcal{A}_{n_{2}}^{\sharp}(\det\nolimits^{\kappa}\rho_{2}\boxtimes\det\nolimits^{\nu}\tau_{2})$
Condition (A) corresponds to Case (I) in Ibukiyama [15] for Siegel modular forms and aligns with the differential operators constructed for several vector-valued cases in Browning [3]. This type of operator is also treated for Hermitian automorphic forms on $U(p,q)$ in the work of Eischen–Liu [8] (see also [6], [7]).
A representation-theoretic interpretation of $\mathbb{D}$ in more general settings was provided in [34], while Case (II) of [15] for Hermitian modular forms was studied by Dunn [5]. The differential operator $\mathbb{D}$ for Hermitian modular forms was explicitly computed in certain cases in [33], following the methods developed by Ibukiyama [14, 16]. In what follows, we recall the relevant results from [34] and [33]. In those references, only the case $\nu=0$ is treated; however, the results extend easily to the case where $\nu$ is allowed to be non-zero.
[…]
Definition 4.1.
If a polynomial $f(X,Y)\in L_{n,\kappa}$ satisfies
$\sum_{s=1}^{\kappa}\frac{\partial^{2}f}{\partial X_{is}\partial Y_{j,s}}=0\text{ for any }i,j\in\{1,\ldots,n\},$
we say that $f(X,Y)$ is pluriharmonic polynomial.
Proposition 4.2 (Corollary 3.21 and Proposition 3.22 of [34]).
Let $n_{1},n_{2}$ be positive integers with $n_{1}\geq n_{2}\geq 1$, and set $n=n_{1}+n_{2}$. Let $(\bm{k}_{s},\bm{l}_{s})=(\bm{k}_{s,v},\bm{l}_{s,v})_{v}$ be a family of pairs of dominant integral weights satisfying $\ell(\bm{k}_{s,v})\leq n_{s},\ell(\bm{l}_{s,v})\leq n_{s},\ell(\bm{k}_{s,v})+\ell(\bm{l}_{s,v})\leq\kappa_{v}+\nu_{v}$ for $s=1,2$ and all $v$. Let $P_{v}(T)$ be a $\bigl(V_{n_{1},\bm{k}_{1,v},\bm{l}_{1,v}}\otimes V_{n_{2},\bm{k}_{2,v},\bm{l}_{2,v}}\bigr)$-valued polynomial on the space $\mathrm{M}_{n}$ of $n\times n$ matrices, for each $v$. Define the differential operator
[…]
  1. (1)
    The differential operator $\mathbb{D}$ satisfies Condition (A) for ${\det}^{\kappa}\boxtimes{\det}^{\nu}$ and $({\det}^{\kappa}\rho_{n_{1},\bm{k}_{1}}\boxtimes{\det}^{\nu}\rho_{n_{1},\bm{l}_{1}})\boxtimes({\det}^{\kappa}\rho_{n_{2},\bm{k}_{2}}\boxtimes{\det}^{\nu}\rho_{n_{2},\bm{l}_{2}})$ if and only if, for each $v$, the polynomial $P_{v}(T)$ satisfies the following conditions.
    1. (a)
      Define
      $\widetilde{P_{v}}(X_{1},X_{2},Y_{1},Y_{2}):=P_{v}\!\left(\begin{pmatrix}X_{1}\,{}^{t}Y_{1}&X_{1}\,{}^{t}Y_{2}\\ X_{2}\,{}^{t}Y_{1}&X_{2}\,{}^{t}Y_{2}\end{pmatrix}\right)$,
      where $X_{i},Y_{i}\in\mathrm{M}_{n_{i},\kappa_{v}+\nu_{v}}$. Then $\widetilde{P_{v}}$ is pluriharmonic in each pair of variables $(X_{i},Y_{i})$.
    2. (b)
      For $(A_{i},B_{i})\in\mathfrak{K}_{n_{i},v}^{\mathbb{C}}:=\mathrm{GL}_{n_{i}}(\mathbb{C})\times\mathrm{GL}_{n_{i}}(\mathbb{C})$, we have
      $P_{v}\!\left(\begin{pmatrix}A_{1}&0\\ 0&A_{2}\end{pmatrix}T\begin{pmatrix}{}^{t}B_{1}&0\\ 0&{}^{t}B_{2}\end{pmatrix}\right)=\bigl(\rho_{n_{1},(\bm{k}_{1,v},\bm{l}_{1,v})}(A_{1},B_{1})\otimes\rho_{n_{2},(\bm{k}_{2,v},\bm{l}_{2,v})}(A_{2},B_{2})\bigr)P_{v}(T).$
  2. (2)
    There exists a differential operator $\mathbb{D}$ satisfying Condition (A) for ${\det}^{\kappa}\boxtimes{\det}^{\nu}$ and $({\det}^{\kappa}\rho_{n_{1},\bm{k}_{1}}\boxtimes{\det}^{\nu}\rho_{n_{1},\bm{l}_{1}})\boxtimes({\det}^{\kappa}\rho_{n_{2},\bm{k}_{2}}\boxtimes{\det}^{\nu}\rho_{n_{2},\bm{l}_{2}})$ if and only if $\bm{k}_{1}=\bm{l}_{2}$ and $\bm{l}_{1}=\bm{k}_{2}$. If such an operator exists, it is unique up to scalar multiplication.
Remark 4.3.
In particular, the differential operators do not depend on the individual choices of $\kappa$ and $\nu$, but only on their sum $\kappa+\nu$.
Remark 4.4.
The explicit form of the polynomial $P_{v}(T)$ is given in [33] for certain specific cases.
4.2. Hermitian Eisenstein series
We introduce the Hermitian Eisenstein series according to Shimura [31, §16.5]. We fix tuples $\kappa=(\kappa_{v})_{v\in\mathbf{a}},\,\nu=(\nu_{v})_{v\in\mathbf{a}}$ of non-negative integers.
Consider the following subgroups of $G_{n}$ for $r\leq n$:
$L_{n,r}$ $=\left\{\left.\begin{pmatrix}A&0&0&0\\ 0&I_{r}&0&0\\ 0&0&{A^{*}}^{-1}&0\\ 0&0&0&I_{r}\\ \end{pmatrix}\in G_{n}\right|A\in\mathrm{GL}_{n-r}(K)\right\},$
$U_{n,r}$ $=\left\{\begin{pmatrix}I_{n-r}&*&*&*\\ 0&I_{r}&*&0\\ 0&0&I_{n-r}&0\\ 0&0&*&I_{r}\\ \end{pmatrix}\in G_{n}\right\},$
$G_{n,r}$ $=\left\{\begin{pmatrix}I_{n-r}&0&0&0\\ 0&*&0&*\\ 0&0&I_{n-r}&0\\ 0&*&0&*\\ \end{pmatrix}\in G_{n}\right\}.$
Then the subgroups $P_{n,r}=G_{n,r}L_{n,r}U_{n,r}$ are the standard parabolic subgroups of $G_{n}$ and there are natural embeddings $a_{n}:F\hookrightarrow A^{+}_{n}$, $t_{n,r}:\mathrm{GL}_{n-r}(K)\hookrightarrow L_{n,r}$ and $s_{n,r}:G_{r}\hookrightarrow G_{n,r}$. Define $G_{n,r,v},G_{n,r}(\mathbb{A}),L_{n,r,v},L_{n,r}(\mathbb{A})$, etc. in the same way as $G_{n,v},G_{n}(\mathbb{A})$, etc.
By the Iwasawa decomposition, $G_{n}(\mathbb{A})$ (resp. $G_{n,v}$) can be decomposed as $G_{n}(\mathbb{A})=P_{n,r,\mathbb{A}}\mathfrak{K}_{n,\mathbb{A}}$ (resp. $G_{n,v}=P_{n,r,v}\mathfrak{K}_{n,v}$).
We consider an unramified Hecke character $\chi$ of $K$, and whose infinite component $\chi_{\infty}$ is given by
$\chi_{\infty}(x)=\prod_{v\in\mathbf{a}}|x_{v}|^{-\kappa_{v}+\nu_{v}}\,x_{v}^{\kappa_{v}-\nu_{v}}.$ (4.1)
For each place $v$ of $F$, we put
$\chi_{v}=\prod_{w\mid v}\chi_{w}.$
Fix such a Hecke character $\chi$.
we put
$\delta(g)=\det\!\left(C\bm{i}_{n}+D\right)$
for $g=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in\mathrm{U}_{n,n}.$
Definition 4.5.
Let $s\in\mathbb{C}$. For $g=t_{n,0}(A)\mu k\in G_{n,v}$ with $A\in\mathrm{GL}_{n}(K_{v})$, $\mu\in U_{n,0,v}$, and $k\in\tilde{K}_{n,v}$, define
$\varepsilon_{n,\kappa,\nu,v}(g,s;\chi)=\begin{cases}|\det({A^{*}}A)|_{v}^{\,s}\,\chi_{v}(\det A^{*})&(v\in\mathbf{h}),\\[6.0pt] \det(g)^{\nu_{v}}|\delta(g)|^{\,\kappa_{v}+\nu_{v}-2s}\,\delta(g)^{-\kappa_{v}-\nu_{v}}&(v\in\mathbf{a}).\end{cases}$
Set
$\varepsilon_{n,\kappa,\nu}(g,s;\chi)=\prod_{v}\varepsilon_{n,\kappa,\nu,v}(g_{v},s;\chi).$
The Hermitian Eisenstein series on $G_{n}(\mathbb{A})$ is defined by
$E_{n,\kappa,\nu}(g,s;\chi)=\sum_{\gamma\in P_{n,0}\backslash G_{n}}\varepsilon_{n,\kappa,\nu}(\gamma g,s;\chi).$
For $\theta\in\mathfrak{K}_{n,0}$, we further define
$E_{n,\kappa,\nu}^{\theta}(g,s;\chi)=E_{n,\kappa,\nu}(g\theta^{-1},s;\chi).$
The Hermitian Eisenstein series $E_{n,\kappa,\nu}(g,s;\chi)$ and $E^{\theta}_{n,\kappa,\nu}(g,s;\chi)$ converge absolutely and locally uniformly for $\mathrm{Re}(s)>n$ (see, for example, Shimura [30]).
Proposition 4.6 ([31, Proposition 17.7]).
Let $\mu$ be a positive integer such that $\mu\geq n$. If $\kappa_{v}+\nu_{v}=\mu$ for any $v\in\mathbf{a}$, then $E_{n,\mu}(g,\mu/2;\chi)=E_{n,\kappa,\nu}(g,s;\chi)$ belongs to $\mathcal{A}^{\sharp}_{n}({\det}^{\kappa}\boxtimes{\det}^{\nu})$, except when $\mu=n+1$, $F=\mathbb{Q}$, $\chi=\chi_{K/F}^{n+1}$, where $\chi_{K/F}$ is the quadratic character associated to quadratic extension $K/F$.
Let $\bm{k},\bm{l}$ be families of dominant integral weights such that $\ell(\bm{k}_{v}),\ell(\bm{l}_{v})\leq r\leq n$, and $\ell(\bm{k}_{v})+\ell(\bm{l}_{v})\leq\kappa_{v}+\nu_{v}$ for each infinite place $v$ of $F$. We put $\rho_{r}={\det}^{\kappa}\rho_{r,\bm{k}}\boxtimes{\det}^{\nu}\rho_{r,\bm{l}}$ and $\rho_{n}={\det}^{\kappa}\rho_{n,\bm{k}}\boxtimes{\det}^{\nu}\rho_{n,\bm{l}}$.
Definition 4.7.
We define
$\varepsilon(f)_{\kappa,\nu,v}^{n}(g,s;\chi)=\left\{\begin{array}[]{ll}\left\lvert\det\left({A_{r}^{*}}A_{r}\right)\right\rvert_{v}^{s}\chi_{v}(\det A_{r}^{*})f(h_{r})&(v\in\mathbf{h}),\\ \left\lvert\delta(g)\delta(h_{r})^{-1}\right\rvert^{\kappa_{v}+\nu_{v}-2s}\rho_{n}(M(g))^{-1}\rho_{r}(M(h_{r}))f(h_{r})\qquad&(v\in\mathbf{a})\end{array}\right$.
for $f\in\mathcal{S}_{r}(\rho_{r})$ ($r<n$) and $g=t_{n,r}(A_{r})\,\mu_{r}\,s_{n,r}(h_{r})\,k\in G_{n,v}$ with $A_{r}\in\mathrm{GL}_{n-r}(K_{v})$, $\mu_{r}\in U_{n,r}$, $h_{r}\in G_{r,v}$ and $k\in\mathfrak{K}_{n,v}$. Then, we put
$\varepsilon(f)_{\kappa,\nu}^{n}(g,s;\chi)=\prod_{v}\varepsilon(f)_{\kappa,\nu,v}^{n}(g_{v},s;\chi)$
and define the Hermitian Klingen-Eisenstein series $\left[f\right]_{\kappa,\nu}^{n}(g,s;\chi)$ on $G_{n}(\mathbb{A})$ associated with $f$ by
$\left[f\right]_{\kappa,\nu}^{n}(g,s;\chi)=\sum_{\gamma\in P_{n,r}\backslash G_{n}}\varepsilon(f)_{\kappa,\nu}^{n}(\gamma g,s;\chi).$
4.3. Pullback formula
Let $n_{1}$ and $n_{2}$ be positive integers with $n_{1}\geq n_{2}$, and put $n=n_{1}+n_{2}$. Let $m=m_{F}$ be the number of infinite places of $F$. Let $\kappa=(\kappa_{v})_{v},\nu=(\nu_{v})_{v}$ be families of positive integers, such that $\kappa_{v}+\nu_{v}=\mu$ for each infinite place $v$ of $F$, where $\mu$ is the fixed positive integer as above. and let $\bm{k}=(\bm{k}_{v})_{v}$ and $\bm{l}=(\bm{l}_{v})_{v}$ be families of dominant integral weights such that $\ell(\bm{k}_{v})<n_{2}$, $\ell(\bm{l}_{v})\leq n_{2}$, and $\ell(\bm{k}_{v})+\ell(\bm{l}_{v})\leq\kappa_{v}+\nu_{v}$ for each infinite place $v$ of $F$. We put $\rho_{n_{1}}={\det}^{\kappa}\rho_{n_{1},\bm{k}}\boxtimes{\det}^{\nu}\rho_{n_{1},\bm{l}}$ and $\rho^{\prime}_{n_{2}}={\det}^{\kappa}\rho_{n_{2},\bm{l}}\boxtimes{\det}^{\nu}\rho_{n_{2},\bm{k}}$, and fix the differential operator $\mathbb{D}_{\bm{k},\bm{l}}=\mathbb{D}_{\bm{k},\bm{l}}^{(n_{1},n_{2})}=P_{\bm{k},\bm{l}}(\pi^{+})$ satisfying condition (A) for ${\det}^{\kappa}\boxtimes{\det}^{\nu}$ and $\rho_{n_{1}}\boxtimes\rho^{\prime}_{n_{2}}$, as defined in [33, Theorem 5.13]. Hereafter, $\mathbb{D}_{\bm{k},\bm{l}}$ denotes this uniquely determined differential operator.
Let $\chi_{K/F}$ be the quadratic character associated with the quadratic extension $K/F$. For a Hecke eigenform $f$ on $\widetilde{G}^{+}_{n}(\mathbb{A})$ of weight $(\rho,V)$, and a Hecke character $\eta$ of $K$, we set
$D(s,f;\eta)=L(s-n+\tfrac{1}{2},\,f\otimes\eta,\,\mathrm{St})\cdot\left(\prod_{i=0}^{2n-1}L_{F}(2s-i,\,\eta\cdot\chi_{K/F}^{i})\right)^{-1},$
where $L(*,f\otimes\eta,\mathrm{St})$ denotes the standard $L$-function of degree $2n$ attached to $f\otimes\eta$ as an automorphic form on $G_{n}(\mathbb{A})$. Here, $L_{F}(*,\eta)$ (resp. $L_{F}(*,\eta\cdot\chi_{K/F})$) is the Hecke $L$-function attached to $\eta$ (resp. $\eta\cdot\chi_{K/F}$).
We put
$g^{\natural}=\begin{pmatrix}0&I_{n}\\ I_{n}&0\end{pmatrix}g\begin{pmatrix}0&I_{n}\\ I_{n}&0\end{pmatrix},\qquad f^{\dagger}_{\chi}(g)=\chi(g)\,\overline{f(g^{\natural})},$
for a Hecke character $\chi$ of $K$ satisfying (4.1). Then, if $f\in\mathcal{A}_{n}^{\sharp}(\det^{\kappa}\rho\boxtimes\det^{\nu}\tau)$, we have $f^{\dagger}_{\chi}\in\mathcal{A}_{r}^{\sharp}(\det^{\kappa}\tau\boxtimes\det^{\nu}\rho)$.
Theorem 4.8.
Assume that $\mathrm{Re}(s)>n$. Let the above notation and assumptions stand. Then, for any Hecke cusp form $f\in\mathcal{S}^{\sharp}_{n_{2}}(\rho^{\prime}_{n_{2}})$, we have
$\left(f,\,(\mathbb{D}_{\bm{k},\bm{l}}E_{n,\kappa})(\iota_{n_{1},n_{2}}(g_{1},*),\overline{s};\chi)\right)=c(s,\rho^{\prime}_{n_{2}})\,D(s,f;\overline{\chi})\,\overline{[f^{\dagger}_{\chi}]_{\kappa}^{n_{1}}(g_{1},s;\chi)},$
where $c(s,\rho^{\prime}_{n_{2}})$ is a function depending only on $\rho^{\prime}_{n_{2}}$, which is defined in [34].
Proof.
The case where $\nu=0$ is treated in [34, Theorem 5.10], and the general case can be proved by a similar argument. ∎
The constant $c(s,\rho^{\prime}_{n_{2}})$ in the theorem is generally not explicitly stated, but in certain cases its value can be computed. We set $|\bm{k}|=\sum_{v,i}k_{v,i}$ and $|\bm{l}|=\sum_{v,i}l_{v,i}$ for the tuple of dominant integral weights $(\bm{k},\bm{l})$.
Proposition 4.9.
Let the above notions and assumptions stand. Assume that $\kappa_{v}+\nu_{v}=\mu$ for all $v\in\mathbf{a}$. Then the constant $c(\mu/2,\rho^{\prime}_{n_{2}})$ appearing in Theorem 4.8 is given by
$c(\mu/2,\rho^{\prime}_{n_{2}})={}$ $2^{-2mn_{2}^{2}-m(n_{2}-2)\mu+|\bm{k}|+|\bm{l}|}\,\pi^{mn_{2}^{2}}$
$\cdot\prod_{v\in\mathbf{a}}\Biggl(\frac{\prod_{i=1}^{\ell(\bm{k}_{v})}\bigl(k_{v,i}+\ell(\bm{k}_{v})-i\bigr)!}{\prod_{1\leq i<j\leq\ell(\bm{k}_{v})}\bigl(k_{v,i}-k_{v,j}+j-i\bigr)}\cdot\frac{\prod_{i=1}^{\ell(\bm{l}_{v})}\bigl(l_{v,i}+\ell(\bm{l}_{v})-i\bigr)!}{\prod_{1\leq i<j\leq\ell(\bm{l}_{v})}\bigl(l_{v,i}-l_{v,j}+j-i\bigr)}$
$\hskip 71.13188pt\cdot\prod_{i=1}^{n_{2}}\frac{1}{\bigl(k_{v,i}+l_{v,i}+\mu-i\bigr)_{(n_{2})}}\Biggr),$
where $(x)_{(r)}=x(x-1)\cdots(x-r+1)$ denotes the descending Pochhammer symbol.
Proof.
The case where $\nu=0$ is treated in [33, Theorem 6.6], and the general case can be proved by a similar argument. ∎
5. Integrality of Hermitian automorphic forms
Hereafter, we restrict our attention to the case where $F=\mathbb{Q}$, $K=\mathbb{Q}(\sqrt{-D_{K}})$ with discriminant $D_{K}$ $(D_{K}>0)$. In this setting, finite places are represented by prime numbers $p$, while there is a unique infinite place, which we denote by $\infty$. We also denote $\chi_{K/F}$ by $\chi_{-D_{K}}$.
Since
$\left|\mathbb{Q}^{\times}\backslash\mathbb{A}^{\times}/\prod_{p}\mathbb{Z}_{p}^{\times}\mathbb{R}^{\times}\right|=1$
in (2.2), we write the complete set of representatives in (2.2) as $\mathfrak{G}_{0}=\{\gamma_{11},\ldots,\gamma_{1h_{2}}\}=\{\gamma_{1},\ldots,\gamma_{h}\}$.
Lemma 5.1.
There is an isomorphism
$\widetilde{\mathcal{A}}_{n}(\rho,\mathbf{1})\;\cong\;\mathcal{A}_{n}(\rho),\qquad f\;\mapsto\;f|_{G_{n}(\mathbb{A})},$
where $\mathbf{1}$ denotes the trivial character of $\mathbb{R}_{>0}$.
Proof.
We set
[…]
$=G_{n}(\mathbb{Q})\cap\bigl(\gamma_{s}\mathfrak{K}_{n,0}\gamma_{s}^{-1}\cdot G_{n,\infty}\bigr)$
for $\gamma_{s}\in\mathfrak{G}_{0}$. Then, by (2.2), we have
[…]
$\;\cong\;\mathbb{R}_{>0}\times\bigl(G_{n}(\mathbb{Q})\backslash G_{n}(\mathbb{A})/\mathfrak{K}_{n,0}\bigr),$
and the assertion follows. ∎
Henceforth, we identify these spaces via the above isomorphism.
In the following, we further assume that $\mu$ is a positive even integer, $\kappa=\nu=\mu/2$ and a Hecke character $\chi$ is trivial. For simplicity, we adopt the following notation:
$E_{n,\mu}(g):=E_{n,\kappa,\nu}(g,\mu/2;\chi),\qquad f^{\dagger}:=f^{\dagger}_{\chi},\qquad[f]_{\mu}^{n}(g):=[f]_{\kappa,\nu}^{n}(g,\mu/2;\chi),$
for any Hecke eigenform $f\in\mathcal{A}^{\sharp}_{r}(\rho)$.
We realize the representation space of $\rho_{n,(\bm{k},\bm{l})}$ as $\mathbb{C}[U,V]_{(\bm{k},\bm{l})}$, using bideterminants, and define $R[U,V]_{(\bm{k},\bm{l})}$ for a commutative ring $R$ in the same manner.
For a subring $R\subset\mathbb{C}$, we set
$\mathcal{A}^{\sharp}_{n}(\rho_{n})(R)=\{\,f\in\mathcal{A}^{\sharp}_{n}(\rho_{n})\mid A_{f}(\gamma_{j},S)\in R[U,V]_{(\bm{k},\bm{l})},\ \forall\,\gamma_{j},\ S\in\mathrm{Her}_{n}(K)\,\}.$
and
$\mathcal{S}^{\sharp}_{n}(\rho_{n})(R)=\mathcal{A}^{\sharp}_{n}(\rho_{n})(R)\cap\mathcal{S}_{n}(\rho_{n}).$
To study the integrality properties of Eisenstein series, we recall results of Shimura [30, 29] on the Fourier expansions of Eisenstein series.
Let $b_{p}(s;S)$ denote the local Siegel series defined in [30, Section 18] for $p\in\mathbf{h}$, $s\in\mathbb{C}$ and $S\in\mathrm{Her}(K_{p})_{\geq 0}$. For $S\in\mathrm{Her}(K_{p})_{\geq 0}$ with $\det S\neq 0$, there exists a polynomial $F_{p}(X;S)$ in $X$ with constant term $1$ and coefficients in $\mathbb{Z}$ such that
$b_{p}(s;S)=\prod_{i=1}^{\lfloor(n+1)/2\rfloor}\bigl(1-p^{-s-2i}\bigr)\prod_{i=1}^{\lfloor n/2\rfloor}\bigl(1-\xi_{p}p^{-s-2i+1}\bigr)\,F_{p}\bigl(p^{-s};S\bigr),$ (5.1)
where
[…]
Remark 5.2.
The Siegel series in the inert case is treated in [12]. The split case is discussed in [27].
Proposition 5.3 ([30, Propositions 18.14 and 19.2], [29]).
Let $\mu$ be a positive even integer with $\mu\geq n$ and $\mu\neq n+1$. Then, for any non-degenerate $S\in\mathrm{Her}_{n}(K)$, the Fourier coefficient
$A_{n,\mu}(g,S):=A_{E_{n,\mu}(g)}(g,S)$
of the Hermitian Eisenstein series $E_{n,\mu}(g)$ is given by
$A_{n,\mu}(\mathbf{m}(A),S)=$ $\;\epsilon_{n}\cdot 2^{n}(D_{K}^{\lfloor{n/2}\rfloor}\det S)^{\mu-n}\prod_{p\in\mathbf{h}}|\det(A_{p}A_{p}^{*})|_{p}^{\,n-\mu/2}\cdot|\det(A_{\infty}A_{\infty}^{*})|^{\,\mu/2}$
$\quad\cdot\prod_{i=0}^{n-1}L(1-\mu+i,\chi_{-D_{K}}^{i})^{-1}\cdot\prod_{p\in\mathbf{c}}F_{p}(p^{-\mu};S[A_{p}])\cdot\exp(-\operatorname{Tr}(2\pi SA_{\infty}{A_{\infty}^{*}})),$
where $\epsilon_{n}=\begin{cases}-1&(n\equiv 2\pmod{4})\\ 1&(\text{otherwise})\end{cases}$, for $A\in\mathrm{GL}_{n}(\mathbb{A}_{K})$. Here, $\mathbf{c}$ is the subset of $\mathbf{h}$ consisting of all the finite places $p$ of the following two types: (i) $p$ is ramified in $K$, (ii) $p$ is unramified in $K$ and $\det({A_{p}^{*}}SA_{p})\not\in\mathbb{Z}_{p}^{\times}$.
In particular, we have
$A_{n,\mu}(\gamma_{i},S)=$ $\;\epsilon_{n}\cdot 2^{n}(D_{K}^{\lfloor{n/2}\rfloor}\det S)^{\mu-n}\,\prod_{i=0}^{n-1}L(1-\mu+i,\chi_{-D_{K}}^{i})^{-1}$
$\qquad\cdot\chi(z_{i})^{-1}\prod_{p\in\mathbf{h}}|\det(z_{i,p}{z_{i,p}^{*}})|_{p}^{\,n-\mu/2}\prod_{p\in\mathbf{c}}F_{p}\bigl(p^{-\mu};S[\mathbf{t}(z_{i,p})]\bigr)$
for $\gamma_{i}\in\mathfrak{G}_{0}$.
Here, the local factors $F_{p}(X;S)$ are known to satisfy the following functional equation (cf. Ikeda [18]):
$F_{p}(p^{-2n}X^{-1};S)=\chi_{-D_{K}}((-1)^{n/2}\det S)^{s_{n}}(p^{n}X)^{-v_{v}(D_{K}^{\lfloor{n/2}\rfloor}\det S)}F_{p}(X;S),$
where $D_{K/F}$ is the discriminant ideal of $K/F$ and $s_{n}=0$ if n is odd, $s_{n}=1$ if n is even. Using this functional equation, we have
$A_{n,\mu}(\mathbf{m}(A),S)=$ $2^{n}\prod_{i=0}^{n-1}L(1-\mu+i,\chi_{-D_{K}}^{i})^{-1}|\det(AA^{*})|_{\mathbb{A}}^{\mu/2}$
$\qquad\cdot\prod_{p\in\mathbf{c}}F_{p}(p^{\mu-2n};{A_{p}^{*}}SA_{p})\exp(-\operatorname{Tr}(2\pi SA_{\infty}{A_{\infty}^{*}})).$ (5.2)
Lemma 5.4.
If $S\in\mathrm{Her}_{n}(\mathcal{O}_{K})_{>0}$ is a non-degenerate Hermitian matrix, then $F_{p}(p^{-n}X;S)\in\mathbb{Z}[X]$. In particular, if $\mu\geq n$, then $F_{p}(p^{\mu-2n};S)\in\mathbb{Z}$.
Proof.
This follows easily from Lemma 4.2.3 of Katsurada [21]. ∎
We put
$\widetilde{E}_{n,\mu}(g)=\left(\prod_{i=0}^{n-1}L(1-\mu+i,\chi_{-D_{K}}^{\,i})\right)E_{n,\mu}(g)=\left(\prod_{i=0}^{n-1}-\frac{B_{\mu-i,\chi_{-D_{K}}^{\,i}}}{\mu-i}\right)E_{n,\mu}(g),$
where $B_{m,\chi}$ denotes the generalized Bernoulli number attached to a Dirichlet character $\chi$. By Proposition 5.3, (5), and Lemma 5.4, we obtain the following result.
Proposition 5.5.
Let $\mu$ be an integer, and assume that either $\mu>n+2$ or $\mu=n$. Then the Eisenstein series $E_{n,\mu}(g)$ and $\widetilde{E}_{n,\mu}(g)$ belong to $\mathcal{A}_{n}({\det}^{\mu})(\mathbb{Q})$, and the denominators of their Fourier coefficients are bounded.
Moreover, for a prime $p$ satisfying the following conditions, we have $\widetilde{E}_{n,\mu}(g)\in\mathcal{A}_{n}({\det}^{\mu})(\mathbb{Z}_{(p)})$:
  • $p>\mu+1$;
  • $p$ does not divide $D_{K}$;
  • either $\mu\geq 2n$, or $\gamma_{i,p}\in\mathrm{GL}_{n}(\mathcal{O}_{K_{p}}^{\times})$ for all $1\leq i\leq h$.
Remark 5.6.
By the von Staudt-Clausen theorem, if $p>\mu+1$, then $p$ does not divide the denominator of $B_{\mu-i}$ for any $0\leq i\leq n-1$. On the other hand, if a prime $p$ does not divide $D_{K}$, then the generalized Bernoulli number $B_{m,\chi_{D}}$ is a $p$-integer for every integer $m$ (see Leopoldt [35]).
As before, we assume that the dominant integral weights $\bm{k}$ and $\bm{l}$ satisfy $\ell(\bm{k})+\ell(\bm{l})\leq\mu$, and we set $\ell(\bm{k},\bm{l})=\max\{\ell(\bm{k}),\ell(\bm{l})\}$. For an integer $r\geq\ell(\bm{k},\bm{l})$, we define $\rho_{r}={\det}^{\mu/2}\rho_{r,\bm{k}}\boxtimes{\det}^{\mu/2}\rho_{r,\bm{l}}$ and $\rho^{\prime}_{r}={\det}^{\mu/2}\rho_{r,\bm{l}}\boxtimes{\det}^{\mu/2}\rho_{r,\bm{k}}$, and set $|\rho_{r}|=|\bm{k}|+|\bm{l}|+r\mu$. For an element $v\otimes v^{\prime}\in V_{\rho_{r}}\otimes V_{\rho^{\prime}_{r}}$, we often omit the symbol $\otimes$ and write $vv^{\prime}=v\otimes v^{\prime}$ as a notation.
In order to present the pullback formula for Eisenstein series in a convenient form, we introduce the following notation.
We set
$\mathbb{L}_{F}(m,\mu)$ $=\prod_{i=0}^{m-1}L_{F}(1-\mu+i,\chi_{-D_{K}}^{i}),$
$\mathbb{L}(s,f)$ $=D_{K}^{r/2}\cdot\prod_{j=1}^{r}\Gamma_{\mathbb{C}}\!\left(s+\tfrac{\mu}{2}+k_{j}-j+\tfrac{1}{2}\right)\Gamma_{\mathbb{C}}\!\left(s+\tfrac{\mu}{2}+l_{j}+r-j-\tfrac{1}{2}\right)\frac{L(s,f,\mathrm{St})}{(f,f)},$
$\mathcal{C}_{n,\mu}(f)$ $=\frac{\mathbb{L}_{F}(n,\mu)}{\mathbb{L}_{F}(2r,\mu)}\,\mathbb{L}\bigl((\mu+1)/2-r,f\bigr),$
where $\Gamma_{\mathbb{C}}(s)=2(2\pi)^{-s}\Gamma(s)$, and $f\in\mathcal{S}_{r}^{\sharp}(\rho_{r})$ is a Hecke cusp form.
We remark that, for
$f=(f_{1},\ldots,f_{h})^{\sharp}\in\mathcal{A}_{r}^{\sharp}(\rho_{r}),\qquad f_{i}(Z)=\sum_{T}a\bigl(f_{i},T\bigr)\,\exp\!\bigl(2\pi\sqrt{-1}\,\operatorname{Tr}(TZ)\bigr),$
a straightforward computation shows that one can write
$f^{\dagger}=(f_{1}^{\dagger},\ldots,f_{h}^{\dagger})^{\sharp}\in\mathcal{A}_{r}^{\sharp}(\rho^{\prime}_{r}),$
where the Fourier coefficients are given by
$a(f_{i}^{\dagger},T)=(\sqrt{-1})^{|\rho_{r}|}\,\overline{a(f_{i},T)}.$
This description relies on the assumption that the level is $1$. If the level is not $1$, such a simple expression in terms of Fourier coefficients is no longer available.
Let $n_{1}$ and $n_{2}$ be positive integers with $n_{1}\geq n_{2}$, and put $n=n_{1}+n_{2}$. We define
$\mathcal{E}_{\mu}(g_{1},g_{2}):=\mathcal{E}^{(n_{1},n_{2})}_{\mu,\bm{k},\bm{l}}(g_{1},g_{2})=(2\pi\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\,\mathbb{D}_{\bm{k},\bm{l}}\,\widetilde{E}_{n_{1}+n_{2},\mu}\bigl(\iota_{n_{1},n_{2}}(g_{1},g_{2})\bigr).$
Remark 5.7.
The normalized Eisenstein series satisfies
$\widetilde{E}_{n,\mu}(g)=\mathbb{L}_{F}(n,\mu)\,E_{n,\mu}(g).$
We also define a constant $M_{r}$ by
$M_{r}:=\max\!\left(\left\{\mu+k_{j}-j-r+1,\;\mu+l_{j}-j\right\}_{1\leq j\leq r}\cup\left\{\mu,\;D_{K},\;k_{1}+l_{1}+\mu-1\right\}\right).$
Lemma 5.8.
Let $n_{1}$ and $n_{2}$ be positive integers with $n_{1}\geq n_{2}$, and put $n=n_{1}+n_{2}$. For $r\geq\ell(\bm{k},\bm{l})$, let $\{f^{(r)}_{d}\mid d=1,\ldots,e_{r}\}$ be an orthogonal basis of $\mathcal{S}_{r}^{\sharp}(\rho_{r})$ with respect to the Petersson inner product, consisting of Hecke cusp forms. Assume that $\mu>n+2$ or $\mu=n$. Then we have
$\mathcal{E}_{\mu}(g_{1},g_{2})=\sum_{r=\ell(\bm{k},\bm{l})}^{n_{2}}c_{r}\sum_{d=1}^{e_{r}}\overline{\mathcal{C}_{n,\mu}\bigl(f^{(r)}_{d}\bigr)}\,[f^{(r)}_{d}]_{\mu}^{n_{1}}(g_{1})\,[f^{(r)\dagger}_{d}]_{\mu}^{n_{2}}(g_{2}),$
where $c_{r}$ is a $p$-unit rational number for any prime $p$ satisfying $p>M_{r}$.
Proof.
This lemma is a direct corollary of Theorem 4.8, whose proof follows the same strategy as, for example, [1, Theorem 5.8]. The constant $c_{r}$ is given explicitly by
$c_{r}=(2\pi\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\,c(\mu/2,\rho^{\prime}_{r})\cdot\frac{D_{K}^{-r/2}\prod_{i=0}^{2r-1}L_{F}(1-\mu+i,\chi_{-D_{K}}^{i})\,L_{F}(\mu-i,\chi_{-D_{K}}^{i})^{-1}}{\prod_{j=1}^{r}\Gamma_{\mathbb{C}}(\mu+k_{j}-j-r+1)\,\Gamma_{\mathbb{C}}(\mu+l_{j}-j)}.$
The $p$-unit property of $c_{r}$ follows from Proposition 4.9, Remark 5.6. ∎
We expand $\mathcal{E}_{\mu}(g_{1},g_{2})$ into a Fourier series with respect to the variables $g_{1}$ and $g_{2}$, and write
$\mathcal{E}_{\mu}(g_{1},g_{2})=\sum_{\begin{subarray}{c}S_{1}\in\mathrm{Her}_{n_{1}}(K)\\ S_{2}\in\mathrm{Her}_{n_{2}}(K)\end{subarray}}\epsilon_{\mu,\bm{k},\bm{l}}^{(n_{1},n_{2})}(g_{1},g_{2};S_{1},S_{2})=\sum_{\begin{subarray}{c}S_{1}\in\mathrm{Her}_{n_{1}}(K)\\ S_{2}\in\mathrm{Her}_{n_{2}}(K)\end{subarray}}\epsilon_{\mu}(g_{1},g_{2};S_{1},S_{2}).$
Then we have,
$\epsilon_{\mu}(g_{1},g_{2};S_{1},S_{2})=\sum_{r=\ell(\bm{k},\bm{l})}^{n_{2}}(\sqrt{-1})^{|\rho_{r}|}c_{r}\sum_{d=1}^{e_{r}}\overline{\mathcal{C}_{n,\mu}\bigl(f^{(r)}_{d}\bigr)}\ A_{[f^{(r)}_{d}]_{\mu}^{n_{1}}}(g_{1};S_{1})\,\overline{A_{[f^{(r)}_{d}]_{\mu}^{n_{2}}}(g_{2};S_{2})}.$ (5.3)
We put
$\mathcal{F}_{\mu}(g_{1},\gamma^{(n_{2})};S_{2})$ $=\mathcal{F}^{(n_{1},n_{2})}_{\mu}(g_{1},\gamma^{(n_{2})};S_{2})$
$=(\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\sum_{S_{1}\in\mathrm{Her}_{n_{1}}(K)}\epsilon_{\mu}(g_{1},\gamma^{(n_{2})};S_{1},S_{2})$
$=\sum_{r=\ell(\bm{k},\bm{l})}^{n_{2}}(-1)^{r\mu/2}c_{r}\sum_{d=1}^{e_{r}}\overline{\mathcal{C}_{n,\mu}\bigl(f^{(r)}_{d}\bigr)}\overline{A_{[f^{(r)}_{d}]_{\mu}^{{n_{2}}}}(\gamma^{(n_{2})},S_{2})}[f^{(r)}_{d}]_{\mu}^{{n_{1}}}(g_{1}),$
$\mathcal{G}_{\mu}(\gamma^{(n_{1})},g_{2};S_{1})$ $=\mathcal{G}^{(n_{1},n_{2})}_{\mu}(\gamma^{(n_{1})},g_{2};S_{1})$
$=(\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\sum_{S_{2}\in\mathrm{Her}_{n_{2}}(K)}\epsilon_{\mu}(\gamma^{(n_{1})},g_{2};S_{1},S_{2})$
$=(\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\sum_{r=\ell(\bm{k},\bm{l})}^{n_{2}}c_{r}\sum_{d=1}^{e_{r}}\overline{\mathcal{C}_{n,\mu}\bigl(f^{(r)}_{d}\bigr)}A_{[f^{(r)}_{d}]_{\mu}^{{n_{1}}}}(\gamma^{(n_{1})},S_{1})[f^{(r)\dagger}_{d}]_{\mu}^{{n_{2}}}(g_{2})$
for $\gamma^{(n_{i})}\in\mathfrak{G}^{(n_{i})}_{0}$ and $S_{i}\in\mathrm{Her}_{n_{i}}(K)$ ($i=1,2$).
Proposition 5.9.
We have
$(\sqrt{-1})^{|\rho_{n_{2}}|}\,\epsilon_{\mu}(\gamma_{(n_{1})},\gamma_{(n_{2})};S_{1},S_{2})\in(V_{\rho_{n_{1}}}\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Q})$
for any Hermitian matrices $S_{i}\in\mathrm{Her}_{n_{i}}(K)$ and $\gamma_{(n_{1})}\in\mathfrak{G}^{(n_{1})}_{0}$, $\gamma_{(n_{2})}\in\mathfrak{G}^{(n_{2})}_{0}$.
More precisely,
$(\sqrt{-1})^{|\rho_{n_{2}}|}\,\epsilon_{\mu}(\gamma_{(n_{1})},\gamma_{(n_{2})};S_{1},S_{2})\in(V_{\rho_{n_{1}}}\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Z}_{(p)}),$
if a prime $p$ satisfies the following conditions:
  • $p>\mu+1$.
  • $p$ does not divide $D_{K}$.
  • All coefficients of $P_{\bm{k},\bm{l}}(T)$ are contained in $(V_{\rho_{n_{1}}}\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Z}_{(p)})$.
  • Either $\mu\geq 2n$, or $\gamma_{i,p}\in\mathrm{GL}_{n}(\mathcal{O}_{K_{p}}^{\times})$ for all $1\leq i\leq h$.
In particular, under the above assumptions, we have
$\mathcal{F}_{\mu}(g_{1},\gamma_{(n_{2})};S_{2})\in(\mathcal{A}_{n_{1}}^{\sharp}(\rho_{n_{1}})\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Z}_{(p)}),\quad\mathcal{G}_{\mu}(\gamma_{(n_{1})},g_{2};S_{1})\in(V_{\rho_{n_{1}}}\otimes\mathcal{A}_{n_{2}}^{\sharp}(\rho^{\prime}_{n_{2}}))(\mathbb{Z}_{(p)}).$
Proof.
Since for any $f=(f_{1},\ldots,f_{h})^{\sharp}\in\mathcal{A}^{\sharp}_{n}({\det}^{\mu})$, we have
$2^{-(|\bm{k}|+|\bm{l}|)}\,\mathbb{D}_{\bm{k},\bm{l}}f=(P_{\bm{k},\bm{l}}(\partial_{Z})f_{1},\ldots,P_{\bm{k},\bm{l}}(\partial_{Z})f_{h})^{\sharp},$
where $\partial_{Z}=(\partial/\partial Z_{ij})$. Noting that $P$ is a homogeneous polynomial of degree $|\bm{k}|+|\bm{l}|$, this implies that the action of $(2\pi\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\mathbb{D}_{\bm{k},\bm{l}}$ preserves the rationality properties of $f$. (See also [1, Proposition 6.6].) ∎
The following lemma is almost obvious from the definition.
Lemma 5.10.
Let $F_{1},\ldots,F_{s}$ be linearly independent elements of $\mathcal{A}^{\sharp}_{n}(\rho_{n})$ consisting of Hecke eigenforms. Then there exist
$C=(C_{v})\in\prod_{v\in\mathbf{a}}\mathrm{M}_{\ell(\bm{k}_{v}),n}(\mathbb{Z}),\qquad D=(D_{v})\in\prod_{v\in\mathbf{a}}\mathrm{M}_{\ell(\bm{l}_{v}),n}(\mathbb{Z}),$
elements $S_{1},\dots,S_{s}\in\mathrm{Her}_{r}(K)$, and $h_{1},\dots,h_{s}\in G_{r}(\mathbb{A}_{0})$ such that
$\Delta:=\det\bigl(A_{F_{i}}(h_{j},S_{j})\bigr)_{1\leq i,j\leq s}\neq 0.$
Remark 5.11.
The elements $h_{1},\ldots,h_{r}$ in the lemma might be chosen from $\mathfrak{G}_{0}$.
Theorem 5.12.
Let $\mu$ be a positive even integer with $\mu\geq 2n$. If $F$ is a Hecke eigenform in $\mathcal{A}^{\sharp}_{n}(\rho_{n})$, then
$F\in\mathcal{A}^{\sharp}_{n}(\rho_{n})(\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{C}.$
In particular,
$\mathcal{A}^{\sharp}_{n}(\rho_{n})=\mathcal{A}^{\sharp}_{n}(\rho_{n})(\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{C}.$
Proof.
This proof is based on an idea suggested by Professor Katsurada. Let $\iota_{n,n}\colon G_{n}\times G_{n}\to G_{2n}$ be the diagonal embedding. Using Lemma 2.5, we expand $\mathcal{E}_{\mu}(g_{1},g_{2})$ as a Fourier series with respect to $g_{2}$, and write
$\mathcal{E}_{\mu}(g_{1},\gamma_{j})=\sum_{S\in{\Lambda_{n_{2}}}_{\geq 0}}\mathcal{F}_{\mu}(g_{1},\gamma_{j};S).$
Since Proposition 5.5 holds and $(2\pi\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}\mathbb{D}_{\bm{k},\bm{l}}$ preserves rationality, it follows that
$\mathcal{F}_{\mu}(g_{1},\gamma_{j};S)(U_{2},V_{2})\in\mathcal{A}^{\sharp}_{n_{1}}(\rho^{\prime}_{n_{1}})(\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Q},$
when viewed as a function of $g_{1}$. Although $\mathcal{F}_{\mu}$ may be regarded as a polynomial in $U_{1},V_{1},U_{2},V_{2}$, we write $\mathcal{F}_{\mu}$ to emphasize that it is a polynomial in $U_{2},V_{2}$.
Applying Lemma 5.10 for $\{F_{1}\ldots,F_{s}\}=\left\{[f^{(r)\dagger}_{d}]_{\mu}^{n}\mid m\leq r\leq n,1\leq d\leq e_{r}\right\}\subset\mathcal{A}^{\sharp}_{n}(\rho_{n})$, we take $C$,$D$, $S_{j}$ and $h_{j}$ such that $\Delta\neq 0$. We may assume that $F=F_{1}=[f^{(r)}_{d}]_{\mu}^{n}$. Then, we have
$\frac{c(f^{(r)}_{d})}{(f^{(r)}_{d},f^{(r)}_{d})}F(g_{1})$
$=\Delta^{-1}\begin{vmatrix}\mathcal{F}_{\mu}(g_{1},h_{1};S_{1})(C,D)&A_{F_{2}}(h_{1},S_{1})(C,D)&\cdots&A_{F_{s}}(h_{1},S_{1})(C,D)\\ \vdots&\vdots&\ddots&\vdots\\ \mathcal{F}_{\mu}(g_{1},h_{d};S_{d})(C,D)&A_{F_{2}}(h_{d},S_{d})(C,D)&\cdots&A_{F_{s}}(h_{d},S_{d})(C,D)\end{vmatrix}$
Since $c(f^{(r)}_{d})$ is not zero, We have $F(g_{1})\in\mathcal{A}^{\sharp}_{n}(\rho_{n})(\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{C}$. ∎
For a Hecke eigenform $f$ and $T\in\mathcal{H}_{n}$, we denote by $\lambda_{f}(T)$ the Hecke eigenvalue of $T$ with respect to $f$, that is,
$T\cdot f=\lambda_{f}(T)\,f.$
We define the field $\mathbb{Q}(f)$ generated over $\mathbb{Q}$ by all Hecke eigenvalues of $f$ by
$\mathbb{Q}(f)=\mathbb{Q}(\{\lambda_{f}(T)\mid T\in\mathcal{H}_{n}\}),$
and call it the Hecke field of $f$. It is well known that $\mathbb{Q}(f)$ is a CM field.
For $f\in\mathcal{A}_{r}^{\sharp}(\rho_{r})$ and $\sigma\in\mathrm{Aut}(\mathbb{C})$, we define $f^{\sigma}:=(f_{1}^{\sigma},\ldots,f_{h}^{\sigma})^{\sharp}$, where each component is given by
$f_{i}^{\sigma}(Z)=\sum_{T\in{\Lambda_{r}}_{\geq 0}}\sigma(a(f_{i},T))\exp(2\pi\sqrt{-1}\,\operatorname{Tr}(TZ)).$
Here, $\sigma$ acts on $\mathbb{C}[U,V]_{(\bm{k},\bm{l})}$ by applying $\sigma$ to each coefficient.
The rationality of the spaces of automorphic forms and of Hermitian Klingen-Eisenstein series can be obtained by the same method as Appendix A of Mizumoto [22].
Proposition 5.13.
Let $n\geq r$ be a positive integer such that $\mu\geq n+r+2$ or $\mu=n+r$. Let $0\neq f\in\mathcal{S}^{\sharp}_{r}(\rho_{r})(\mathbb{Q}(f))$ be a Hecke cusp form. Then the following assertions hold:
  1. (1)
    There exists an orthogonal basis $\{f_{1},\ldots,f_{d}\}$ of $\mathcal{S}^{\sharp}_{r}(\rho_{r})$ consisting of Hecke cusp forms such that $f_{j}\in\mathcal{S}^{\sharp}_{r}(\rho_{r})(\mathbb{Q}(f_{j}))$ for each $j$.
  2. (2)
    We have
    $\left(\mathcal{C}_{n,\mu}(f)\right)^{\sigma}=\mathcal{C}_{n,\mu}(f^{\sigma})\quad\text{for all }\sigma\in\mathrm{Aut}(\mathbb{C}).$
    In particular,
    $\mathcal{C}_{n,\mu}(f)\in\mathbb{Q}(f).$
  3. (3)
    We have
    $([f]_{\mu}^{n})^{\sigma}=[f^{\sigma}]_{\mu}^{n}\quad\text{for all }\sigma\in\mathrm{Aut}(\mathbb{C}).$
    In particular,
    $f]_{\mu}^{n}\in\mathcal{A}^{\sharp}_{n}(\rho_{n})(\mathbb{Q}(f)).$
Corollary 5.14.
Let $n$ be a positive integer, and let $(\bm{k},\bm{l})$ be dominant integral weights such that $k_{n}+l_{n}\geq 2n$. For any Hecke cusp form $f\in\mathcal{S}^{\sharp}_{n}\bigl(\rho_{n,(\bm{k},\bm{l})}\bigr)$ and any integer $s_{0}$ satisfying $0\leq s_{0}\leq(k_{n}+l_{n})/2-n$, we have
$\frac{L\bigl(s_{0}+\frac{1}{2},\,f,\,\mathrm{St}\bigr)}{D_{K}^{n/2}\cdot\pi^{|\bm{k}|+|\bm{l}|+2n(s_{0}+1/2)-n}\cdot(f,f)}\in\mathbb{Q}(f).$
Remark 5.15.
More precise results on the algebraicity of special values of standard $L$-functions for scalar-valued Hermitian modular forms were obtained by Shimura [31]. See also Bouganis [2].
Remark 5.16.
When $n=1$ and the class number of $K$ is $1$, we have
$L\bigl(s_{0}+\tfrac{1}{2},\,f^{\sharp},\,\mathrm{St}\bigr)=L\bigl(s_{0}+\tfrac{k}{2},\,f\bigr)L\bigl(s_{0}+\tfrac{k}{2},\,f\otimes\chi_{K}\bigr),$
where $f\in S_{k}(\mathrm{SL}_{2}(\mathbb{Z}))$ is a classical Hecke cusp form and $L(s,f)$ denotes its Dirichlet $L$-series:
$L(s,f)=\sum_{n=1}^{\infty}a(n,f)n^{-s}=\prod_{p}(1-a(p,f)p^{-s}+p^{k-1-2s})^{-1}.$
Moreover, the Petersson norm $(f^{\sharp},f^{\sharp})$ on the Hermitian upper half-plane coincides with the classical one (cf. [17, Appendix to §10]). Thus, our result is consistent with the period relations shown by Shimura [28].
We set $\mathcal{H}_{n}=\otimes_{p}^{\!\prime}\mathcal{H}_{n,p}$ to be the (global) Hecke algebra, and put $\mathcal{H}_{n}(\mathbb{Q})=\mathcal{H}_{n}\otimes_{\mathbb{Z}}\mathbb{Q}$.
We say that an element $T\in\mathcal{H}_{n}(\mathbb{Q})$ is integral with respect to $\mathcal{A}^{\sharp}_{n}(\rho)(\mathbb{Z})$ if
$T\cdot f\in\mathcal{A}^{\sharp}_{n}(\rho)(\mathbb{Z})\quad\text{for all }f\in\mathcal{A}^{\sharp}_{n}(\rho)(\mathbb{Z}).$
We denote by $\mathcal{H}_{n}^{(\bm{k},\bm{l})}$ the subset of $\mathcal{H}_{n}(\mathbb{Q})$ consisting of all elements that are integral with respect to $\mathcal{A}^{\sharp}_{n}(\rho_{n,(\bm{k},\bm{l})})$.
From (3.1),(3.1),(3.2) and (3.3), we see that
Proposition 5.17.
Under the notation of 3.3, if $k_{n}+l_{n}\geq 2n$, then
$\underset{p:\text{inert}}{\bigotimes\nolimits^{\prime}}\mathcal{H}_{n,p}\subset\mathcal{H}_{n}^{(\bm{k},\bm{l})}.$
Moreover, if we additionally assume that the class number of $K$ is $1$, then
$\mathcal{H}_{n}\subset\mathcal{H}_{n}^{(\bm{k},\bm{l})}.$
Lemma 5.18.
Suppose that $k_{n}\geq 2n$. Let $f\in\mathcal{S}^{\sharp}_{n}(\rho_{n,(\bm{k},\bm{l})})$ be a Hecke eigenform. Then $\lambda_{f}(T)$ belongs to $\mathcal{O}_{\mathbb{Q}(f)}$ for any $T\in\mathcal{H}_{n}^{(\bm{k},\bm{l})}$.
Proof.
This follows immediately from Theorem 5.12 and the definition of $\mathcal{H}_{n}^{(\bm{k},\bm{l})}$. ∎
Now we define the integral ideal $\mathfrak{A}(f)$ of $\mathbb{Q}(f)$ attached to a Hecke cusp form $f\in\mathcal{S}^{\sharp}_{n}(\rho_{n})(\mathbb{Q}(f))$, and state the following integrality lemma, formulated along the same lines as in Mizumoto [25, 24, 23].
Put $V=\bigoplus_{\tau}\mathbb{C}f^{\tau}$, where $\tau$ runs over all embeddings of $\mathbb{Q}(f)$ into $\mathbb{C}$, and let $V^{\perp}$ be the orthogonal complement of $V$ in $\mathcal{S}^{\sharp}_{n}(\rho_{n})$. Let $\nu(f)$ (resp. $\kappa(f)$) denote the exponent of the finite abelian group $\mathcal{S}^{\sharp}_{n}(\rho_{n})(\mathbb{Z})/(V(\mathbb{Z})\oplus V^{\perp}(\mathbb{Z}))$ (resp. $\mathcal{O}_{\mathbb{Q}(f)}/\mathbb{Z}[\lambda_{f}(T)\mid T\in\mathcal{H}_{n}^{(\bm{k},\bm{l})}]$). We define
$\mathfrak{A}(f)=\kappa(f)\nu(f)\mathfrak{d}(\mathbb{Q}(f)),$
where $\mathfrak{d}(\mathbb{Q}(f))$ denotes the different of $\mathbb{Q}(f)/\mathbb{Q}$.
If $f\in\mathcal{S}^{\sharp}_{n}(\rho_{n})(\mathbb{Q}(f))$ has a Fourier coefficient $A_{f}(\gamma_{j},S)$ with $\gamma_{j}\in\mathfrak{G}_{0}$ and $S\in\mathrm{Her}_{n}(K)_{\geq 0}$ such that $v_{\mathfrak{p}}(A_{f}(\gamma_{j},S))=0$ for any prime ideal $\mathfrak{p}$ of $\mathbb{Q}(f)$, we say that $f$ has a primitive Fourier coefficient.
Lemma 5.19 (Integrality lemma).
Assume that $\mu\geq 2n$. Let $f\in\mathcal{S}^{\sharp}_{n}(\rho_{n})(\mathbb{Q}(f))$ be a Hecke cusp form having a primitive Fourier coefficient, and let $L$ be an algebraic number field. Then for any $g\in S_{k}(\Gamma_{n})(\mathcal{O}_{L})$, we have
$\frac{(f,g)}{(f,f)}\in\mathfrak{A}(f)^{-1}\cdot\mathcal{O}_{L\cdot\mathbb{Q}(f)}.$
Proof.
The proof is identical to that of [24, Theorem 4.1] and [25, Lemma 3.1]. ∎
For non-negative integers $\nu$ and $m$ with $\ell(\bm{k},\bm{l})\leq m<\nu$, let $e_{(\bm{k},\bm{l})}(\nu,m)$ denote the exponent of the finite abelian group $\mathcal{S}^{\sharp}_{m}(\rho_{m})(\mathbb{Z})]_{\mu}^{\nu}/[\mathcal{S}^{\sharp}_{m}(\rho_{m})]_{\mu}^{\nu}(\mathbb{Z})$. We agree that $e_{(\bm{k},\bm{l})}(\nu,m)=1$ if $\mathcal{S}^{\sharp}_{m}(\rho_{m})=\{0\}$, and we put
$\eta_{(\bm{k},\bm{l})}(\nu):=\prod_{m=\ell(\bm{k},\bm{l})}^{\nu-1}e_{(\bm{k},\bm{l})}(\nu,m).$
Proposition 5.20.
Assume that $\mu\geq 2n$. Let $f\in\mathcal{S}^{\sharp}_{r}(\rho_{r})(\mathbb{Q}(f))\ (\ell(\bm{k},\bm{l})\leq r\leq n)$ be a Hecke cusp form having a primitive Fourier coefficient. If a prime $p$ satisfies the following conditions:
  • $p>\mu+1$,
  • $p$ does not divide $D_{K}$,
  • all coefficients of $P_{\bm{k},\bm{l}}(T)$ are contained in $(V_{\rho_{n_{1}}}\otimes V_{\rho^{\prime}_{n_{2}}})(\mathbb{Z}_{(p)})$,
  • either $\mu\geq 2(n+r)$, or $\gamma_{i,p}\in\mathrm{GL}_{n}(\mathcal{O}_{K_{p}}^{\times})$ for all $1\leq i\leq h$,
then we have
$c_{r}\overline{\mathcal{C}_{n+r,\mu}(f)}A_{[f]_{\mu}^{{n}}}(\gamma,S)\in V_{\rho_{n}}\bigl(\eta_{(\bm{k},\bm{l})}(r)^{-1}\mathfrak{A}(f)^{-1}\mathbb{Z}_{(p)}\bigr)$
for any $\gamma\in\mathfrak{G}_{0}$ and $S\in\mathrm{Her}_{n}(K)_{\geq 0}$.
Proof.
We consider $\mathrm{pr}(\mathcal{G}^{(n,r)}_{\mu}(\gamma,g_{2};S))$, where $\mathrm{pr}:\mathcal{A}^{\sharp}_{n}(\rho_{n})\to\mathcal{S}^{\sharp}_{n}(\rho_{n})$ denotes the projection onto the space of cusp forms. We have
$\mathrm{pr}(\mathcal{G}^{(n,r)}_{\mu}(\gamma,g_{2};S))=(\sqrt{-1})^{-(|\bm{k}|+|\bm{l}|)}c_{r}\sum_{d=1}^{e_{r}}\overline{\mathcal{C}_{n+r,\mu}\bigl(f^{(r)}_{d}\bigr)}A_{[f^{(r)}_{d}]_{\mu}^{{n}}}(\gamma,S)f^{(r)\dagger}_{d}(g_{2}),$
and by Proposition 5.9, $\mathrm{pr}(\mathcal{G}^{(n,r)}_{\mu}(\gamma,g_{2};S))$ belongs to $(V_{\rho_{n}}\otimes\mathcal{A}_{r}^{\sharp}(\rho^{\prime}_{r}))(\eta_{(\bm{k},\bm{l})}(r)^{-1}\mathbb{Z}_{(p)})$ (cf. [24, Proposition 6.2]). Applying Lemma 5.19 to $f^{\dagger}$ and $\mathrm{pr}(\mathcal{G}^{(n,r)}_{\mu}(\gamma,g_{2};S))$, we obtain the desired result. ∎
6. Main theorem
The notation and assumptions in the previous sections remain in force.
Definition 6.1.
Let $F,G\in\mathcal{A}^{\sharp}_{n}(\rho_{n,(\bm{k},\bm{l})})$ be Hecke eigenforms with $k_{n}\geq 2n$, and let $\mathfrak{p}$ be a prime ideal of $\mathbb{Q}(F,G)=\mathbb{Q}(F)\cdot\mathbb{Q}(G)$. If
$\lambda_{F}(T)\equiv\lambda_{G}(T)\pmod{\mathfrak{p}}\quad\text{for all }T\in\mathcal{H}^{(\bm{k},\bm{l})}_{n}$,
we say that $F$ and $G$ are Hecke congruent, and we denote this by
$F\equiv_{ev}G\pmod{\mathfrak{p}}$.
For $r\geq\ell(\bm{k},\bm{l})$, let $\{f^{(r)}_{1},\ldots,f^{(r)}_{e_{r}}\}$ be an orthogonal basis of $\mathcal{S}_{r}^{\sharp}(\rho_{r})$ with respect to the Petersson inner product, consisting of Hecke cusp forms and satisfying the following conditions:
  • Each $f^{(r)}_{d}$ belongs to $\mathcal{S}_{r}^{\sharp}(\rho_{r})(\mathbb{Q}(f^{(r)}_{d}))$.
  • The basis is permuted under the action of $\mathrm{Aut}(\mathbb{C})$.
  • Each $f^{(r)}_{d}$ has a primitive Fourier coefficient.
The existence of such a basis follows from Proposition 5.13.
We are now in a position to state our main result.
Theorem 6.2.
Let $n$ be a positive integer, let $\mu$ be a positive even integer with $\mu\geq 2n$, and let $(\bm{k},\bm{l})$ be a pair of dominant integral weights such that $\ell(\bm{k}),\ell(\bm{l})\leq n$. Let $f\in\mathcal{S}_{r}^{\sharp}(\rho_{r})(\mathbb{Q}(f))$ be a Hermitian automorphic form which is a constant multiple of $f^{(r)}_{1}$ for some integer $r$ satisfying $\ell(\bm{k},\bm{l})\leq r\leq n$.
Assume that a prime ideal $\mathfrak{p}$ of $\mathbb{Q}(f)$, together with $S_{0}\in\mathrm{Her}_{n}(K)_{>0}$ and $\gamma_{0}\in\mathfrak{G}_{0}$, satisfies the following conditions:
  1. (1)
    We have
    $v_{\mathfrak{p}}\!\left(\overline{\mathcal{C}_{2n,\mu}\bigl(f\bigr)}A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})\overline{A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})}\right)=:-\alpha<0.$
    Note that the above quantity remains unchanged if we replace $f$ by $cf$ for any constant $c\in\mathbb{Q}(f)^{\times}$.
  2. (2)
    If $\ell(\bm{k},\bm{l})\leq\nu\leq n-1$, $1\leq d\leq e_{\nu}$, and $(\nu,d)\neq(r,1)$, then
    $v_{\mathfrak{q}}\!\left(\mathcal{C}_{n+\nu,\mu}(f^{(\nu)}_{d})\right)\leq 0$
    for any prime ideal $\mathfrak{q}$ of $\mathbb{Q}(f,f^{(\nu)}_{d})$ lying above $\mathfrak{p}$.
  3. (3)
    The ideals $\mathfrak{A}(f^{(\nu)}_{d})$ are coprime to $\mathfrak{p}$ in $\mathbb{Q}(f,f^{(\nu)}_{d})$ for all $\ell(\bm{k},\bm{l})\leq\nu\leq n-1$ and $1\leq d\leq e_{\nu}$.
  4. (4)
    The rational prime $p$ underlying $\mathfrak{p}$ satisfies the following conditions:
    1. (a)
      $p>\mu+1$;
    2. (b)
      $p$ does not divide $D_{K}$;
    3. (c)
      $p$ does not divide $\eta_{(\bm{k},\bm{l})}(n-1)$;
    4. (d)
      all coefficients of $P_{\bm{k},\bm{l}}(T)$ are contained in $(V_{\rho^{\prime}_{n}}\otimes V_{\rho_{n}})(\mathbb{Z}_{(p)})$;
    5. (e)
      either $\mu\geq 4n$, or $\gamma_{i,p}\in\mathrm{GL}_{n}(\mathcal{O}_{K_{p}}^{\times})$ for all $1\leq i\leq h$;
    6. (f)
      $c_{\nu}$ is $p$-unit for all $\ell(\bm{k},\bm{l})\leq\nu\leq n-1$.
Then there exists a Hecke cusp form $F\in\mathcal{S}^{\sharp}_{n}(\rho_{n})$ such that
$F\equiv_{ev}[f]_{\mu}^{n}\pmod{\mathfrak{P}}$
for some prime ideal $\mathfrak{P}$ of $\mathbb{Q}(f,F)$ lying above $\mathfrak{p}$.
Moreover, if $\mathfrak{p}$ is coprime to $\mathfrak{A}(f^{(n)}_{d})$ for all $1\leq d\leq e_{n}$, then there exists a Hecke cusp form $G\in\mathcal{S}^{\sharp}_{n}(\rho_{n})$ such that
$G\equiv_{ev}[f]_{\mu}^{n}\pmod{\mathfrak{P}^{\prime\alpha}}$
for some prime ideal $\mathfrak{P}^{\prime}$ of $\mathbb{Q}(f,G)$ lying above $\mathfrak{p}$.
Before proving Theorem 6.2, we prepare the following lemmata.
Lemma 6.3.
Assume the same hypotheses as in Theorem 6.2. Let $\ell(\bm{k},\bm{l})\leq\nu\leq n-1$ and $1\leq d\leq e_{\nu}$, and assume that $(\nu,d)\neq(r,1)$. Then the Fourier coefficients of
$\overline{\mathcal{C}_{2n,\mu}\bigl(f^{(\nu)}_{d}\bigr)}\,\overline{A_{[f^{(\nu)}_{d}]_{\mu}^{n}}(\gamma_{0},S_{0})}\,[f^{(\nu)}_{d}]_{\mu}^{n_{1}}(g_{1})$
are $\mathfrak{p}$-integral; that is, they are $\mathfrak{q}$-integral elements of $V_{\rho_{n}}\otimes V_{\rho^{\prime}_{n}}\bigl(\mathbb{Q}(f,f^{(\nu)}_{d})\bigr)$ for every prime ideal $\mathfrak{q}$ of $\mathbb{Q}(f,f^{(\nu)}_{d})$ lying above $\mathfrak{p}$.
Proof.
For any $\gamma\in\mathfrak{G}_{0}$ and $S\in\mathrm{Her}_{n}(K)_{\geq 0}$, assumptions (3) and (4c) together with Proposition 5.20 imply that
$c_{\nu}\overline{\mathcal{C}_{2n,\mu}\bigl(f^{(\nu)}_{d}\bigr)}\,A_{[f^{(\nu)}_{d}]_{\mu}^{n}}(\gamma,S)\,\overline{A_{[f^{(\nu)}_{d}]_{\mu}^{n}}(\gamma_{0},S_{0})}$ $=\frac{\mathbb{L}_{F}(2n,\mu)}{\mathbb{L}_{F}(n+\nu,\mu)}\cdot\frac{v}{c_{\nu}\,\mathcal{C}_{n+\nu,\mu}\bigl(f^{(\nu)}_{d}\bigr)}$
$=\prod_{i=n+\nu}^{2n-1}L_{F}(1-\mu+i,\chi_{-D_{K}}^{i})\cdot\frac{v}{c_{\nu}\,\mathcal{C}_{n+\nu,\mu}\bigl(f^{(\nu)}_{d}\bigr)}$,
where $v$ is a $\mathfrak{p}$-integral element of $V_{\rho_{n}}\otimes V_{\rho^{\prime}_{n}}\bigl(\mathbb{Q}(f,f^{(\nu)}_{d})\bigr)$.
By Remark 5.6 and assumptions (4a) and (4b), each factor $L_{F}(1-\mu+i,\chi_{-D_{K}}^{i})$ is $\mathfrak{p}$-integral for $n+\nu\leq i\leq 2n-1$. Therefore, by assumptions (2) and (4f), the assertion follows. ∎
The key to the proof of the first part of the main theorem is the following lemma, which is proved in the same way as Katsurada [20, Lemma 5.1] (see also [1, Lemma 6.10]).
Lemma 6.4.
Let $F_{1},\ldots,F_{d}$ be Hecke eigenforms in $\mathcal{A}^{\sharp}_{n}(\rho_{n})$ which are linearly independent over $\mathbb{C}$. Let $K=\mathbb{Q}(F_{1},\ldots,F_{d})$ be the composite field of their Hecke fields, $\mathcal{O}$ its ring of integers, and let $\mathfrak{p}$ be a prime ideal of $K$. Let $G(Z)\in(\mathcal{A}^{\sharp}_{n}(\rho_{n})\otimes V_{\rho^{\prime}_{n}})(\mathcal{O}_{(\mathfrak{p})})$, and assume that the following conditions hold:
  1. (1)
    The form $G$ is expressed as $G(Z)=\sum_{i=1}^{d}c_{i}F_{i}(Z)$ with $c_{i}\in V_{\rho^{\prime}_{n}}$.
  2. (2)
    There exist $S\in\mathrm{Her}_{n}(K)_{\geq 0}$ and $\gamma\in\mathfrak{G}_{0}$ such that $c_{1}A_{F_{1}}(\gamma,S)\in(V_{\rho_{n}}\otimes V_{\rho^{\prime}_{n}})(K)$ and $v_{\mathfrak{p}}(c_{1}A_{F_{1}}(\gamma,S))<0$.
Then there exists an index $i\neq 1$ such that $F_{i}\equiv_{ev}F_{1}\mod\mathfrak{p}$.
Proof of Theorem 6.2.
By assumption (4), we have
$\mathcal{F}^{(n,n)}_{\mu}(g_{1},\gamma;S_{2})\in(\mathcal{A}_{n}^{\sharp}(\rho_{n})\otimes V_{\rho^{\prime}_{n}})(\mathbb{Z}_{(p)})$
for any $\gamma\in\mathfrak{G}_{0}$ and $S_{2}\in\mathrm{Her}_{n}(K)_{\geq 0}$. Hence, by Lemma 6.3, we obtain
$c_{r}\,\overline{\mathcal{C}_{2n,\mu}(f)}\,\overline{A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})}\,[f]_{\mu}^{n}(g_{1})+\sum_{d=1}^{e_{n}}\beta_{d}\,f^{(n)}_{d}(g_{1})\equiv 0\pmod{(\mathcal{O}_{\mathbb{Q}(f)})_{(\mathfrak{p})}}$, (6.1)
where
$\beta_{d}=(-1)^{(n-r)\mu/2}\,c_{n}\,\overline{\mathcal{C}_{2n,\mu}(f^{(n)}_{d})}\,\overline{A_{f^{(n)}_{d}}(\gamma_{0},S_{0})}\in V_{\rho^{\prime}_{n}}\bigl(\mathbb{Q}(f^{(n)}_{d})\bigr)$.
Here the congruence $\equiv 0\pmod{(\mathcal{O}_{\mathbb{Q}(f)})_{(\mathfrak{p})}}$ means that all Fourier coefficients of the left-hand side lie in
$(V_{\rho_{n}}\otimes V_{\rho^{\prime}_{n}})\bigl((\mathcal{O}_{\mathbb{Q}(f)})_{(\mathfrak{p})}\bigr).$
As in the proof of Proposition 5.20, we note that
$\sum_{d=1}^{e_{n}}\beta_{d}\,f^{(n)}_{d}(g_{1})\in(\mathcal{A}_{n}^{\sharp}(\rho_{n})\otimes V_{\rho^{\prime}_{n}})(\mathbb{Q}).$
By assumptions (1) and (4f), we have
$v_{\mathfrak{p}}\!\left(c_{r}\,\overline{\mathcal{C}_{2n,\mu}(f)}\,A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})\,\overline{A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})}\right)=-\alpha<0.$ (6.2)
Hence Lemma 6.4 applies and yields the first assertion.
Let $\{v_{1},\ldots,v_{t}\}$ be a fixed basis of $V_{\rho_{n}}$. Write
$A_{f^{(n)}_{d}}(\gamma_{0},S_{0})=\sum_{j=1}^{t}a_{dj}v_{j},\qquad a_{dj}\in\mathbb{Q}(f^{(n)}_{d})$.
By multiplying each $f^{(n)}_{d}$ by a suitable element of $\mathbb{Q}(f^{(n)}_{d})^{\times}$ and renumbering if necessary, we may assume
$a_{d1}=\begin{cases}1&(1\leq d\leq e_{0}),\\ 0&(e_{0}+1\leq d\leq e_{n}),\end{cases}$
and that
$v_{\mathfrak{P}^{\prime}}\!\left(c_{n}\,\overline{\mathcal{C}_{2n,\mu}(f^{(n)}_{1})}\right)\leq-\alpha$ (6.3)
for some prime ideal $\mathfrak{P}^{\prime}$ of $\mathbb{Q}(f,f^{(n)}_{1})$ lying above $\mathfrak{p}$. This is possible since
$v_{\mathfrak{p}}\!\left(\sum_{d=1}^{e_{n}}\beta_{d}\,A_{f^{(n)}_{d}}(\gamma_{0},S_{0})\right)=v_{\mathfrak{p}}\!\left(c_{r}\,\overline{\mathcal{C}_{2n,\mu}(f)}\,A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})\,\overline{A_{[f]_{\mu}^{n}}(\gamma_{0},S_{0})}\right)=-\alpha.$
In particular, (6.1) and (6.2) imply that $e_{0}\geq 1$.
For any $T\in\mathcal{H}_{n}^{(\bm{k},\bm{l})}$, applying the operator $T-\lambda_{[f]_{\mu}^{n}}(T)$ to both sides of (6.1), we obtain
$H(g_{1}):=\sum_{d=1}^{e_{0}}\beta_{d}\,\bigl(\lambda_{f^{(n)}_{d}}(T)-\lambda_{[f]_{\mu}^{n}}(T)\bigr)f^{(n)}_{d}(g_{1})\equiv 0\pmod{(\mathcal{O}_{\mathbb{Q}(f)})_{(\mathfrak{p})}}$.
There exists a $\mathfrak{p}$-unit $u\in\mathbb{Q}(f)^{\times}$ such that
$uH(g_{1})\in\mathcal{S}_{n}^{\sharp}(\rho_{n})(\mathcal{O}_{\mathbb{Q}(f)}).$
Applying Lemma 5.19 to $f^{(n)}_{d}$ and $uH(g_{1})$, we obtain
$uc_{n}\,\overline{\mathcal{C}_{2n,\mu}(f^{(n)}_{d})}\,\bigl(\lambda_{f^{(n)}_{d}}(T)-\lambda_{[f]_{\mu}^{n}}(T)\bigr)\in\mathfrak{A}(f^{(n)}_{d})^{-1}\cdot\mathcal{O}_{\mathbb{Q}(f,f^{(n)}_{d})}$
for each $1\leq d\leq e_{0}$. If $\mathfrak{P}^{\prime}$ is coprime to $\mathfrak{A}(f^{(n)}_{1})$, then by (6.3) we obtain
$\lambda_{f^{(n)}_{1}}(T)\equiv\lambda_{[f]_{\mu}^{n}}(T)\pmod{\mathfrak{P}^{\prime\alpha}}$.
This completes the proof. ∎
By following the proof of Theorem 6.2, we obtain the following weaker result under fewer assumptions.
Corollary 6.5.
Under the same notation as in Theorem 6.2, assume that the conditions (1), (4a), (4b), (4d), (4e), and (4f) are satisfied. Then there exists a Hecke eigenform $F\in\mathcal{A}^{\sharp}_{n}(\rho_{n})$ such that
$F\equiv_{ev}[f]_{\mu}^{n}\pmod{\mathfrak{P}}$
for some prime ideal $\mathfrak{P}$ of $\mathbb{Q}(f,F)$ lying above $\mathfrak{p}$.
7. Example
We treat the case $K=\mathbb{Q}(\sqrt{-3})$, whose discriminant is $D_{K}=3$. For each $i\in\{12,26\}$, let $\Delta_{i}\in S_{i}(\mathrm{SL}_{2}(\mathbb{Z}))=S_{i}(\Gamma_{K}^{(1)})$ be the unique cusp form of weight $i$ and level $1$, normalized so that the Fourier coefficient of $q$ is $1$.
By calculating the Gross–Keating invariants, we obtain the following polynomials (for the definition of $F_{p}(X;S)$, see (5.1)):
$F_{3}(X;I_{3})$ $=1+3^{3}X,$
$F_{3}(X;I_{4})$ $=1+2\cdot 3^{5}X+3^{8}X^{2},$
$F_{3}(X;\mathrm{diag}(1,3))$ $=1+3^{4}X^{2},$
$F_{3}(X;\mathrm{diag}(1,1,3))$ $=1-3^{3}(3-1)X+3^{6}X^{2},$
$F_{3}(X;\mathrm{diag}(1,1,1,3))$ $=1+3^{5}X+3^{9}X^{2}+3^{12}X^{3}.$
The computation of the polynomials associated with these Siegel series is based on results from an unpublished note by Ikeda and Katsurada.
[…]
We identify the representation space of $\det^{4}\mathrm{Sym}^{7}\boxtimes\det^{4}\mathrm{Sym}^{7}$ with the space of homogeneous polynomials of degree $7$ in the variables $u_{1},u_{2}$ and of degree $7$ in the variables $v_{1},v_{2}$. Since $\dim\mathcal{S}_{1}(\det^{11}\boxtimes\det^{11})=1$, we have $\mathfrak{A}(\Delta_{22})=1$. By direct computation, we have
$c_{1}=\frac{2^{9}\cdot 3^{3}}{11^{2}\cdot 13^{2}}$
Using SageMath, we compute the constant $\mathcal{C}_{4,8}(\Delta_{22})$ as follows:
$\mathcal{C}_{4,8}(\Delta_{22})$ $=\frac{\mathbb{L}_{F}(4,8)}{\mathbb{L}_{F}(2,8)}\mathbb{L}(7/2,\Delta_{22})$
$=\zeta(1-6)L(1-5,\chi_{-3})\Gamma_{\mathbb{C}}(14)\Gamma_{\mathbb{C}}(14)\frac{L(14,\Delta_{22})L(14,\Delta_{22}\otimes\chi_{-3})}{\sqrt{3}(\Delta_{22},\Delta_{22})}$
$=-\frac{2^{24}\cdot 41}{3^{15}\cdot 5\cdot 19}.$
Furthermore, by Theorem 5.5 of [33], the differential operator $P_{(7,0),(7,0)}(T)$ for $T=(t_{ij})$ is explicitly given by the polynomial
$P_{(7,0),(7,0)}(T)=\frac{1}{2^{12}\cdot 3^{6}\cdot 5^{3}\cdot 7^{2}\cdot 11\cdot 13\cdot 52140059}\sum_{j=0}^{7}(-1)^{j}\binom{7}{j}\binom{20-j}{7-j}A^{7-j}B^{7-j}C^{j}$
where
$A$ $=(u_{1}t_{13}+u_{2}t_{23}),$
$B$ $=(v_{1}t_{31}+v_{2}t_{32}),$
$C$ $=\sum_{i=1}^{2}\sum_{j=1}^{2}u_{i}v_{j}(t_{ij}t_{33}-t_{i3}t_{3j}).$
By substituting this explicit form of $P_{(7,0),(7,0)}(T)$ into the formula in Proposition 5.3, we evaluate the Fourier coefficient of $\Delta_{22}]_{8}^{2}$ at $T_{1}=\begin{pmatrix}2&1\\ 1&2\end{pmatrix}$. A direct computation then yields
$c_{1}\mathcal{C}_{4,8}(\Delta_{22})A_{[\Delta_{22}]_{8}^{2}}(I_{2};T_{1})$
$=$ $-L(1-5,\chi_{-3})\epsilon_{22}(I_{2},I_{1};T_{1},1)$
$=$ $\frac{1}{N}\bigg[-3\cdot 107\cdot 7333\cdot 9623\left(u_{1}^{7}v_{1}^{7}+u_{2}^{7}v_{2}^{7}\right)$
$\qquad+2^{3}\cdot 3\cdot 5^{4}\cdot 7\cdot 11\cdot 17\cdot 22350173\left(u_{1}^{7}v_{1}^{4}v_{2}^{3}+u_{1}^{4}u_{2}^{3}v_{1}^{7}+u_{1}^{3}u_{2}^{4}v_{2}^{7}+u_{2}^{7}v_{1}^{3}v_{2}^{4}\right)$
$\qquad-2^{3}\cdot 7\cdot 17\cdot 6257\cdot 20696549\left(u_{1}^{7}v_{1}v_{2}^{6}+u_{1}^{6}u_{2}v_{2}^{7}+u_{1}u_{2}^{6}v_{1}^{7}+u_{2}^{7}v_{1}^{6}v_{2}\right)$
$\qquad-2^{3}\cdot 3\cdot 7^{4}\cdot 17\cdot 5647\cdot 24113\left(u_{1}^{6}u_{2}v_{1}^{6}v_{2}+u_{1}u_{2}^{6}v_{1}v_{2}^{6}\right)$
$\qquad+5^{2}\cdot 7^{2}\cdot 49277\cdot 22350173\left(u_{1}^{6}u_{2}v_{1}^{3}v_{2}^{4}+u_{1}^{4}u_{2}^{3}v_{1}v_{2}^{6}+u_{1}^{3}u_{2}^{4}v_{1}^{6}v_{2}+u_{1}u_{2}^{6}v_{1}^{4}v_{2}^{3}\right)$
$\qquad-2^{2}\cdot 3\cdot 7^{2}\cdot 324908052197\left(u_{1}^{5}u_{2}^{2}v_{1}^{5}v_{2}^{2}+u_{1}^{2}u_{2}^{5}v_{1}^{2}v_{2}^{5}\right)$
$\qquad+2\cdot 3^{3}\cdot 7\cdot 19\cdot 16603\cdot 22350173\left(u_{1}^{5}u_{2}^{2}v_{1}^{2}v_{2}^{5}+u_{1}^{2}u_{2}^{5}v_{1}^{5}v_{2}^{2}\right)$
$\qquad-2^{4}\cdot 5\cdot 7^{2}\cdot 131\cdot 709\cdot 15756253\left(u_{1}^{4}u_{2}^{3}v_{1}^{4}v_{2}^{3}+u_{1}^{3}u_{2}^{4}v_{1}^{3}v_{2}^{4}\right)\bigg],$
where $N=2^{9}\cdot 3^{7}\cdot 5^{3}\cdot 7^{3}\cdot 11\cdot 13\cdot 52140059$.
Consequently, for a prime $p=41$, we have
$v_{41}\!\left(\overline{\mathcal{C}_{4,22}(\Delta_{22})}\,A_{[\Delta_{22}]_{8}^{2}}(I_{2},T_{1})\cdot\overline{A_{[\Delta_{22}]_{8}^{2}}(I_{2},T_{1})}\right)=-1.$
Since these primes satisfy the hypotheses of Theorem 6.2, we conclude that for each such $p$, there exists a Hecke cusp form $F\in\mathcal{S}^{\sharp}_{2}\bigl(\det^{4}\mathrm{Sym}^{7}\boxtimes\det^{4}\mathrm{Sym}^{7}\bigr)$ such that
$F\equiv[\Delta_{22}]_{8}^{2}\pmod{\mathfrak{P}}$
for some prime ideal $\mathfrak{P}$ of the coefficient field $\mathbb{Q}(F)$ lying above $41$.
[…]
This is the scalar-valued case, and the computation proceeds similarly:
$c_{1}$ $=\frac{2^{7}\cdot 3^{11}\cdot 5}{13},$
$\mathcal{C}_{6,12}(\Delta_{12})$ $=\frac{2^{19}\cdot 7^{2}\cdot 13\cdot 809}{3^{15}\cdot 691}.$
Again by Proposition 5.3,
$c_{1}\,\mathcal{C}_{6,12}(\Delta_{12})\,A_{[\Delta_{12}]_{12}^{3}}(I_{3};T_{2})$ $=\zeta(1-8)L(1-7,\chi_{-3})\,\epsilon_{12}(I_{3},I_{1};T_{2},1)$
$=-\frac{2^{6}\cdot 7^{2}\cdot 79\cdot 89\cdot 57202544699}{3^{3}},$
where $T_{2}=\begin{pmatrix}2&1&0\\ 1&2&0\\ 0&0&1\end{pmatrix}$.
We then obtain
$v_{809}\!\left(\overline{\mathcal{C}_{6,12}(\Delta_{12})}\,A_{[\Delta_{12}]_{12}^{3}}(I_{3},T_{2})\,\overline{A_{[\Delta_{12}]_{12}^{3}}(I_{3},T_{2})}\right)=-1.$
From the above computations together with some additional straightforward verifications, we conclude that $809$ satisfies the assumptions of Corollary 6.5. Hence there exists a Hecke eigenform $F\in\mathcal{A}^{\sharp}_{3}(\det^{6}\boxtimes\det^{6})$ such that
$F\equiv_{ev}[\Delta_{12}]_{12}^{3}\pmod{\mathfrak{P}}$
for some prime ideal $\mathfrak{P}$ of $\mathbb{Q}(F)$ lying above $809$.